### v4

# Some configuration results for area-minimizing cones

Abstract.
We discover some configuration results for area-minimizing cones. In particular, given any closed minimal submanifold in some Euclidean sphere, every cone over the minimal product of sufficiently many copies of the submanifold turns out to be area-minimizing.
Key words and phrases:
Area-minimizing cone, Lawlor criterion
1991 Mathematics Subject Classification:
Primary 28A75, Secondary 53C38
1. Introduction
Let $L$ be a set in the unit sphere $\mathbb{S}^{N}$ in Euclidean space $(\mathbb{R}^{{N+1}},g_{E})$ and $C(L)$ be the cone $\{tx:t\geq 0,\ x\in L\}$ over its link $L$. It is called area-minimizing in geometric measure theory, if the truncated cone $C\bigcap\mathbb{B}^{{N+1}}(1)$ has least mass among all integral currents (see [FF]) with boundary $L$. A fundamental result (Theorem 5.4.3 in [F]) asserts the existence of tangent cone at every point of an area-minimizing rectifiable current and moreover each tangent cone itself area-minimizing. Although ambient Riemannian manifold can be quite arbitrary, after blowing-up those cones always live in standard Euclidean spaces. Lots of efforts have been made in studying important properties and finding examples of them in various situations, e.g., [JS, BdGG, BL, HL, HS, FK, Ch, Law0, Law, NS, XYZ2, z, TZ] and etc.
When $L$ is an embedded closed smooth submanifold, the cone $C(L)$ is singular only at the origin and we say it is a regular cone. In this paper, we search for regular area-minimizing cones and end up with very general configuration results.
Let us get started with a basic question regarding the structure of the category of regular area-minimzing cones.
$(\star)$ Given two regular area-minimizing cones $C(L_{1})$ and $C(L_{2})$ in $\mathbb{R}^{{N_{1}+1}}$ and $\mathbb{R}^{{N_{2}+1}}$ respectively, is the cone over minimal product $L_{1}\times L_{2}$ still area-minimizing in $\mathbb{R}^{{N_{1}+N_{2}+2}}$?
Here the minimal product of $L_{1}$ and $L_{2}$ of dimension $k_{1}$ and $k_{2}$ respectively is
(1.1) $\left(\sqrt{\frac{k_{1}}{{k_{1}+k_{2}}}}L_{1},\sqrt{\frac{k_{2}}{{k_{1}+k_{2}}}}L_{2}\right)\subset\mathbb{S}^{N_{1}+N_{2}+1}.$
Apparently, the minimal product can apply to multiple inputs inductively. Along this line, Lawlor (Theorem 5.1.1, [Law]) completely settled down $(\star)$ when each input $L_{i}$ is $\mathbb{S}^{N_{i}}$, with $i=1,2,\cdots,n$ and $N_{i}\in\mathbb{Z}_{+}$. In particular, whenever $\Sigma_{i=1}^{n}N_{i}>7$, the corresponding cone is area-minimizing. In recent work [TZ], cones of dimension no less than 37 over minimal products of the category of typical minimal submanifolds associated to isoparametric foliations of spheres are confirmed area-minimizing. From these, one can observe that the dimension plays an important role for $(\star)$ and that some minimal submanifolds in spheres themselves cannot span area-minimizing cones but their products do.
In this paper, among others we get the following.
Theorem 1.1.
Given embedded closed minimal submanifolds $\{L_{i}\}$ respectively in Euclidean spheres with $i=1,2,\cdots,n$, every cone over the minimal product $L$ of sufficiently many copies among these $\{L_{i}\}$ is area-minimizing.
The result tells in particular that every closed minimal submanifold with no other assumptions at all can be used as a brick to construct area-minimizing cones in different combinatorial ways. The only payment is that the resulting area-minimizing cone may have high dimension.
Moreover, we establish the following using the terminology from the set theory.
Theorem 1.2.
Regular area-minimizing cones in Euclidean spaces and closed minimal submanifolds in Euclidean spheres share the same cardinality.
Usually, it is not easy to construct regular area-minimizing cones before. But Theorem 1.2 shows that there are a plethora of such kind cones and the cardinality is decided by closed minimal submanifolds in Euclidean spheres.
The main tool is Lawlor’s criterion for a regular cone to be area-minimizing. To be self-contained we briefly explain his method in §2. Some useful information about minimal product will be given in §3. Finally, the proof of Theorems 1.1 and 1.2 and more subtle results (also interesting) by applying Lawlor’s criterion are collected in §4.
Acknowledgment. The author wishes to thank Professor Blaine Lawson for drawing his attention to the powerful work [Law] by Professor Gary Lawlor.
2. Lawlor’s curvature criterion
Area-minimizing hypercone is a key ingredient to solve the famous Bernstein problem [B] about entire minimal graph of codimension one, see [fle, de2, A]. It was Jim Simons who found first non-trivial stable minimal hypercones $C(S^{k}\times S^{k})$ with $k\geq 3$ in [JS] and suspected that they are area-minimizing. In the following year, breakthrough was made in [BdGG] which confirmed those Simons cones’ area-minimality and based on this counterexamples to the Bernstein problem were found over graph domain $\mathbb{R}^{N}$ when $N\geq 8$.
In [HS] a beautiful characterization of area-minimizing hypercones was established. In each side of an area-minimizing hypercone, there is a unique dilation-invariant foliation by minimal hypersurfaces. Therefore, the unit norm vector fields lead to a foliation by their integral curves on $\mathbb{R}^{N}\sim 0$, which is also invariant under dilation. It naturally induces a dual co-degree one calibration form (see [HL]) and anl area-nonincreasing projection (for surfaces of codimension one) to the hypercone.
For a given regular cone, how to determine whether it is area-minimizing or not? The most effective way is to employ the theory of calibration. However, in general, it is not easy to directly find a desired calibration with mild singularity. Lawlor successfully developed a sufficient criterion (sometimes also necessary) looking for area-nonincreasing projection structure in suitable angular neighborhood of the minimal cone instead of in the whole space $\mathbb{R}^{N}$. If the boundary of the neighborhood is mapped to the origin under the projection, then one can send everything outside the neighborhood to the origin. In this way an area-nonincreasing projection can be globally defined. However, its dual calibration has quite “large” singularity.
To understand Lawlor’s criterion, let us recall the following for $k<N$.
Definition 2.1.
Suppose $L^{k}\subset\mathbb{S}^{N}$ is a closed submanifold and $x\in L$. Let $S_{x}$ denote the set of all unit normals of $L$ at $x$. Then the normal radius of $L$ (or $C(L)$) at $x$ is
$R_{x}\triangleq\sup\left\{\,s\in\mathbb{R}_{+}\,:\,\exp_{x}(sv_{x})\not\in L\,\text{ for all }v_{x}\in S_{x}\right\}$
where $\exp_{x}(sv_{x})=(\cos s)x+(\sin s)v_{x}$ is a normal geodsic. The quantity $R(L)\triangleq\inf\{\,R_{x}\,:\,x\in L\,\}$ is called normal radius of $L$ (or $C(L)$).
To establish area-minimizing projection, Lawlor considered foliation generated by rotations of a suitable curve $\gamma_{x}$ and dilations in each normal wedge
$W_{x}\left(\frac{R_{x}}{2}\right)=C\left(\exp_{x}\left(\left[0,\frac{R_{x}}{2}\right]\cdot S_{x}\right)\right).$
[…]
(2.1) $\frac{dr}{d\theta}\leq r\sqrt{r^{2k+2}\cos^{2k}\theta\inf_{x\in L,\,v\in S_{x}}\left(\det\left(\textbf{I}-(\tan\theta)\textbf{h}_{ij}^{v}\right)\right)^{2}-1}$
with $r(0)=\|x\|=1$. Here $\textbf{h}_{ij}^{v}$ means the second fundamental form at $x$ with respect to $v$. To get narrowest curve $\gamma_{x}$, one needs to require that equality holds everywhere in (2.1) and wants to have a smallest $\theta_{0}$ such that $\lim_{\theta\uparrow\theta_{0}}r(\theta)=\infty$. As a result, the boundary of $W_{x}\left(\theta_{0}\right)$ is projected to the origin and global area-decreasing projection gets set up. Such $\theta_{0}$, if existed, is called vanishing angle of $C(L)$. Note that there is no focal point of $L$ in the neighborhood $\bigcup_{x\in L}W_{x}\left(\theta_{0}\right)$ since otherwise that focal point leads to a strictly negative expression under the square root symbol in (2.1).
Lawlor’s curvature criterion is the following.
Theorem 2.2 (Theorem 1.2.1, [Law]).
Given a regular minimal cone $C(L)\subset\mathbb{R}^{N+1}$, if the vanishing angle $\theta_{0}$ exists and $\theta_{0}\leq\frac{R(L)}{2}$, then $C(L)$ is area-minimizing.
By the control
(2.2) $(1-\alpha t)e^{\alpha t}<F(\alpha,t,k+1)\leq\det\left(\textbf{I}-t\,\textbf{h}_{ij}^{v}\right)$
where $t=\tan\theta$, $\alpha=\sup_{x\in L,v\in S_{x}}\|\textbf{h}_{ij}^{v}\|$ and
(2.3) $F(\alpha,t,k+1)=\left(1-\alpha t\sqrt{\frac{k}{k+1}}\right)\left(1+\frac{\alpha t}{\sqrt{k(k+1)}}\right)^{k},$
Lawlor derived two kinds of vanishing angles by employing (2.2) to replace the infimum term in (2.1)
$\theta_{c}(k+1,\alpha)>\theta_{F}(k+1,\alpha)\geq\theta_{0}(C(L)).$
So, if either $\theta_{F}$ or $\theta_{c}$ exists of finite value, then consequently vanishing angle $\theta_{0}$ must exist. For details readers are referred to [Law]. In this paper we will focus on the following nice property.
Lemma 2.3 (Proposition 1.4.2, [Law]).
For $m,r\in\mathbb{Z}_{+}$ and $m>r$, it follows that
(2.4) $\tan\left(\theta_{c}\Big(m,\frac{m}{r}\alpha\Big)\right)<\frac{r}{m}\tan\Big(\theta_{c}(r,\alpha)\Big)\,.$
In fact a ratio type estimate on $\theta_{F}$ can also be established.
Lemma 2.4.
For $m,r\in\mathbb{Z}_{+}$ and $m>r$, it follows that
(2.5) $\tan\left(\theta_{F}\Big(m,\frac{m}{r}\alpha\Big)\right)<\frac{r}{m}\tan\left(\theta_{F}\left(r,\alpha\sqrt{\frac{(m-1)r}{(r-1)m}}\right)\right)\,.$
Although $F$ and $\theta_{F}$ are more accurate and efficient for applying the criterion, however in certain sense Lemma 2.4 is not as good as Lemma 2.3 since the extra factor $\sqrt{\frac{(m-1)r}{(r-1)m}}$ is larger than one. There will be no much difference in the scope of current paper when $m$ is sufficiently large and we save its proof in appendix.
In this paper we will in particular consider the following subclass.
Definition 2.5.
We call a regular minimal cone Type-c (or Type-F) Lawlor cone if it can be verified area-minimizing using the vanishing angle $\theta_{c}$ (or $\theta_{F}$), i.e., $\theta_{c}\leq\frac{R(L)}{2}$ (or $\theta_{F}\leq\frac{R(L)}{2}$).
We will show that the situation in Theorem 1.1 brings us Type-c cones (also Type-F cones). Besides, some considerations regarding $(\star)$ will be made in §4 (and results work for Type-F as well).
3. Minimal product
The (constant) minimal product  [^2] We say constant minimal product to distinguish this from spiral minimal product in [L-Z]. in §1 can naturally induce a nice algorithm to generate minimal submanifold for multiple minimal submanifolds
(3.1) $L_{1}\times\cdots\times L_{n}\triangleq\left(\lambda_{1}L_{1},\cdots,\lambda_{n}L_{n}\right)\subset\mathbb{S}^{N_{1}+\cdots+N_{n}+n-1}.$
where $\lambda_{i}=\sqrt{\frac{k_{i}}{k}}$ with $k_{i}$ the dimension of $L_{i}$ and $k=\sum_{i=1}^{n}k_{i}>0$. In this paper we shall focus on the case that every $k_{i}>0$.
This structure should be a very classical result. There have been many specific situations where minimal product has been employed. For all $L_{i}$ being full spheres, the structure produces classical Clifford type minimal surfaces as already mentioned in §1. When these $L_{i}$ are taken to be R-spaces, [OS] studied whether cones over minimal products of them are area-minimizing. For general minimal submanifolds as inputs, see [X] for a proof through the Takahashi Theorem. In 2016 (arXiv chronology), it was rediscovered independently by [TZ] and [CH] using different methods. Reader may choose either familiar way to understand the structure.
3.1. Estimate on $\alpha^{2}$ for minimal product.
For simplicity, let us fix $n=2$ in the paper and all related results for $n\geq 3$ can be gained by induction. In [TZ], information about the second fundamental form of minimal product has been clearly figured out. Let $k_{1},k_{2}$ be the dimension of $L_{1},L_{2}$ respectively. Take $x_{1}\in L_{1}$ and $x_{2}\in L_{2}$. Together with $\eta_{0}=(\lambda_{2}x_{1},-\lambda_{1}x_{2})$, orthonormal bases $\{\sigma_{1},\cdots,\sigma_{N_{1}-k_{1}}\}$ and $\{\tau_{1},\cdots,\tau_{N_{2}-k_{2}}\}$ of $T_{x_{1}}^{\perp}L_{1}$ and $T_{x_{2}}^{\perp}L_{2}$ lead to an orthonormal basis
$\{(\sigma_{1},0),\cdots,(\sigma_{N_{1}-k_{1}},0),\;(0,\tau_{1}),\cdots,(0,\tau_{N_{2}-k_{2}}),\;\eta_{0}\}$
of the normal space of $L_{1}\times L_{2}$ at point $(\lambda_{1}x_{1},\lambda_{2}x_{2})$ in $\mathbb{S}^{N_{1}+N_{2}+1}$.
Let $A$ stand for the symbol of shape operators. Then with respect to natural tangential orthonormal bases, we have the following relations:
[…]
Let $\alpha_{1}^{2},\,\alpha_{2}^{2},\,\alpha^{2}$ be the supremum norm squares of corresponding second fundamental forms with respect to unit normals for $L_{1},\ L_{2}$ and $L_{1}\times L_{2}$ respectively. Then the following is obvius.
Lemma 3.1.
As marked in the above, it follows that, for $k_{1},k_{2}>0$,
$\alpha^{2}=(k_{1}+k_{2})\max\left\{\frac{\alpha_{1}^{2}}{k_{1}},\,\frac{\alpha_{2}^{2}}{k_{2}},\,1\right\}.$
Proof.
Note that the matrices of shape operators with respect to three subclasses of unit normals $\{(\sigma_{i},0)\}$, $\{(0,\tau_{i})\}$ and $\{\eta_{0}\}$ are mutually perpendicular. ∎
3.2. Normal radius of minimal product
Let $R_{1},\,R_{2},\,R$ be the normal radii of $L_{1},\ L_{2}$ and $L_{1}\times L_{2}$ in corresponding Euclidean spheres respectively.
For $(\lambda_{1}x_{1},\lambda_{2}x_{2})$ in $L_{1}\times L_{2}$, the combination
(3.2) $\nu=a_{1}(\xi_{1},0)+a_{2}(0,\xi_{2})+a_{0}\eta_{0}$
where $\xi_{1}$, $\xi_{2}$ are unit normals of $L_{1}$ and $L_{2}$ at $x_{1}$ and $x_{2}$ respectively, and $a_{1}^{2}+a_{2}^{2}+a_{0}^{2}=1$, can exhaust all unit normals of $L_{1}\times L_{2}$ at $p$. Then, in $\mathbb{S}^{N_{1}+N_{2}+1}$, each unit normal induces a normal geodesic
$\exp_{(\lambda_{1}x_{1},\lambda_{2}x_{2})}(s\nu)$
$=$ $(\cos s)(\lambda_{1}x_{1},\lambda_{2}x_{2})+(\sin s)\nu$
$=$ $\big((\lambda_{1}\cos s+a_{0}\lambda_{2}\sin s)x_{1}+(a_{1}\sin s)\xi_{1},$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ (\lambda_{2}\cos s-a_{0}\lambda_{1}\sin s)x_{2}+(a_{2}\sin s)\xi_{2}\big)$
Therefore, $R$ is the smallest positive number such that there exists
$p\triangleq(\lambda_{1}x_{1},\lambda_{2}x_{2})\in L_{1}\times L_{2}$
satisfying $q\triangleq(\lambda_{1}y_{1},\lambda_{2}y_{2})=\exp_{p}(R\nu_{p})\in L_{1}\times L_{2}$ for some $\nu_{p}\in S_{p}$. Thus,
(3.4) $(\lambda_{1}\cos R+a_{0}\lambda_{2}\sin R)x_{1}+(a_{1}\sin R)\xi_{1}=\lambda_{1}\big((\cos\phi_{1})x_{1}+(\sin\phi_{1})\xi_{1}\big)$
(3.5) $(\lambda_{2}\cos R-a_{0}\lambda_{1}\sin R)x_{2}+(a_{2}\sin R)\xi_{2}=\lambda_{2}\big((\cos\phi_{2})x_{2}+(\sin\phi_{2})\xi_{2}\big)$
for some $\phi_{1},\phi_{2}\in\mathbb{R}$ such that
$y_{1}=(\cos\phi_{1})x_{1}+(\sin\phi_{1})\xi_{1}\text{\ \ \ and \ \ \ }y_{2}=(\cos\phi_{2})x_{2}+(\sin\phi_{2})\xi_{2}.$
Somehow similarly as argued in [TZ], there are three possibilities to realize normal radius $R$, but notice that, a priori, possible values of $\phi_{1}$ and $\phi_{2}$ in the above may not be discrete as in the isoparametric situation in [TZ]. Moreover, except being a point of sphere, the normal radius of a closed submanifold in sphere is no larger than $\pi$.
Case I. $y_{1}\neq x_{1}$ and $y_{2}\neq x_{2}$.
[…]
(3.6) $\cos R=\lambda_{1}^{2}\cos\phi_{1}+\lambda_{2}^{2}\cos\phi_{2}\leq\lambda_{1}^{2}\cos R_{1}+\lambda_{2}^{2}\cos R_{2}$
Case II. $y_{1}\neq x_{1}$ and $y_{2}=x_{2}$ (or the other way around).
Now $\cos\phi_{2}=1$ and (3.6) becomes
(3.7) $\cos R\leq\lambda_{1}^{2}\cos R_{1}+\lambda_{2}^{2}.$
Case III. $y_{1}=x_{1}$ and $y_{2}=x_{2}$.
Since the ambient is Euclidean sphere, if $R$ occurs in this case then $R=2\pi$ (here $2\pi$ corresponds to normal geodesic curve for direction $\nu=\eta_{0}$ when neither $L_{1}$ nor $L_{2}$ contains antipodal points).
So, to search for the value of $R$ one should focus on Case II. Note that it is always true that
$0<R<\min\{\lambda_{1}R_{1},\,\lambda_{2}R_{2}\}<\pi.$
Hence $\sin R\neq 0$. In Case II, following the discussion of [JCX] in fact one can show that no matter what value of $\phi_{1}$ is in (3.4), the triple $(a_{0},a_{1},a_{2})$ can be solved to make (3.4) hold. Namely,
(3.8) $-a_{0}\lambda_{2}=\lambda_{1}\frac{\cos R-\cos\phi_{1}}{\sin R},$
(3.9) $a_{0}\lambda_{1}=\lambda_{2}\frac{\cos R-1}{\sin R},\ \ \$
$a_{2}=0.\ \ \ \ \ \ \ \ \ \ \ \$
These can be read off directly from (3.4) and (3.5). However, the reverse logic will be more important to us, since what we want to solve is this:
Given $\phi_{1}=R_{1}$ for some $x_{1}$ and direction $\xi_{1}$ realizing normal radius $R_{1}$ of $L_{1}$ in $\mathbb{S}^{N_{1}}$, can we get $a_{0},\,a_{1}$ (with $a_{0}^{2}+a_{1}^{2}=1$) and some number $R$ to make (3.4) and (3.5) hold?
If affirmative, we know that the equality of (3.7) stands. This becomes more clear from its solution
(3.10) $\cos R=\lambda_{1}^{2}\cos R_{1}+\lambda_{2}^{2},$
(3.11) $a_{0}\sin R=\lambda_{1}\lambda_{2}(\cos R_{1}-1),$
$a_{1}\sin R=\lambda_{1}\sin R_{1}.\ \ \ \ \ \ \ \ \ \ \ \ \$
automatically with $a_{0}^{2}+a_{1}^{2}=1$.
We remark that the above argument works perfectly for genuine submanifolds of spheres. However, when $L_{1}$ (or $L_{2}$ or both) is a full sphere, the normal radius $R_{1}$ (or $R_{2}$ or both) cannot be well defined as in Definition 2.1. For our purpose, we always define the normal radius of a full sphere $\mathbb{S}^{N_{1}}$ in itself to be $\pi$. The reason is the following. It can be realized as a totally geodesic sphere in $\mathbb{S}^{N_{1}+\ell}$ for $\ell>0$ and now $\mathbb{S}^{N_{1}}$ has normal radius $\pi$. So the above argument works for the minimal product and the last $\ell$ components of $\mathbb{R}^{N_{1}+1+\ell}$ can be completely dropped as they contribute nothing (let $\xi$ be a unit normal vector to the minimal product at some point and if $\xi$ has a nonzero part coming from the last $\ell$ components of $\mathbb{R}^{N_{1}+1+\ell}$ then the geodesic with initial velocity vector $\xi$ in $\mathbb{S}^{N_{1}+\ell+N_{2}}$ will run at least length $\pi$ to return to the minimal product and hence this will not realize the normal radius as the normal radius is strictly less than $\pi$ for the minimal product - not a totally geodesic sphere, actually no more than $\frac{\pi}{2}$ by Lemma 4.7). For example, the normal radius of minimal product $\mathbb{S}^{N_{1}}\times\mathbb{S}^{N_{2}}$ in $\mathbb{S}^{N_{1}+N_{2}+1}$ is realized by $\eta_{0}$ and can reach $\frac{\pi}{2}$ only when $N_{1}=N_{2}$ according to (3.10), also see the beautiful uniform formula in [Law] for normal radius of minimal product among full spheres.
As a result, with the above understanding we arrive at the following.
Lemma 3.2.
Let $R_{1}$ and $R_{2}$ be the normal radii of $L_{1}$ and $L_{2}$ in $\mathbb{S}^{N_{1}}$ and $\mathbb{S}^{N_{2}}$ respectively. Then the normal radius $R\in(0,\pi)$ of $L_{1}\times L_{2}$ in $\mathbb{S}^{N_{1}+N_{2}+1}$ satisfies
(3.12) $\cos R=1-\min\left\{\lambda_{1}^{2}(1-\cos R_{1}),\,\lambda_{2}^{2}(1-\cos R_{2})\right\}.$
Proof.
This is exactly given by the equality situation in Case II and take the larger value of $\lambda_{1}^{2}\cos R_{1}+\lambda_{2}^{2}$ and $\lambda_{1}^{2}+\lambda_{2}^{2}\cos R_{2}$. ∎
Remark 3.3.
Note that the result is in fact valid when $\lambda_{1}$ and $\lambda_{2}$ are just two nonzero real numbers with $a_{1}^{2}+a_{2}^{2}=1$. However, we will restrict ourselves in the minimal product construction unless otherwise specified.
Remark 3.4.
Applying this to $L_{1}\times L_{2}\times L_{3}=(L_{1}\times L_{2})\times L_{3}$, we have
$\cos R(L_{1}\times L_{2}\times L_{3})$
$=$ $1-\min\left\{\frac{k_{1}+k_{2}}{k_{1}+k_{2}+k_{3}}(1-\cos R(L_{1}\times L_{2})),\,\frac{k_{3}}{k_{1}+k_{2}+k_{3}}(1-\cos R_{3})\right\}$
$=$ $1-\min\Bigg\{\frac{k_{1}}{k}(1-\cos R_{1}),\,\frac{k_{2}}{k}(1-\cos R_{2}),\,\frac{k_{3}}{k}(1-\cos R_{3})\Bigg\}.$
By induction, one can get the following formula.
Corollary 3.5.
Let $R_{i}$ be the normal radii of $L_{i}$ in $\mathbb{S}^{N_{i}}$ for $i=1,\cdots,n$. Then the normal radius $R(L_{1}\times\cdots\times L_{n})\in(0,\pi)$ of $L_{1}\times\cdots\times L_{n}$ in $\mathbb{S}^{N_{1}+\cdots+N_{n}+n-1}$ satisfies
(3.13) $\cos R(L_{1}\times\cdots\times L_{n})=1-\min_{1\leq i\leq n}\left\{\lambda_{i}^{2}(1-\cos R_{i})\right\}.$
Remark 3.6.
Not as in §3.1, the formula here does not need $\{L_{i}\}$ to be minimal but depends on the constant multipliers $\{\lambda_{i}\}$ in the construction of the product.
Remark 3.7.
Statement formulated independently in [JCX].
4. Configuration results
In this section we will give a proof of Theorem 1.1 first which applies to arbitrary closed minimal submanifolds in Euclidean spheres. A further glance actually leads us to more results, such as Corollary 4.2 and Theorem 1.2. After that, we will move to some configuration results within the category of Type-c cones.
4.1. For general closed minimal submanifold.
For simplicity, suppose that the set $\{L_{i}\}$ of closed minimal submanifolds in spheres in assumption contains exactly one element $L_{1}^{k_{1}}\subset\mathbb{S}^{N_{1}}$.
Proof of Theorem 1.1. Let $R_{1}>0$ be the normal radius of $L_{1}$ in $\mathbb{S}^{N_{1}}$. Denote the minimal product of $n$-copies of $L_{1}$ by $L_{1}^{\times n}$. Then, by (3.13) and Lemma 4.7 in below, the normal radius $R(L_{1}^{\times n})$ is no more than $\frac{\pi}{2}$ and its exact value is
(4.1) $\arccos\left(1-\frac{1}{n}(1-\cos R_{1})\right)=\sqrt{2(1-\cos R_{1})}\cdot\frac{1}{\sqrt{n}}+\text{ higher order terms }.$
On the other hand, denote by $\alpha_{1}$ and $\alpha$ the supremum norms of second fundamental forms of $L_{1}$ and $L_{1}^{\times n}$ respectively. By induction, Lemma 3.1 implies that
$\alpha=\sqrt{nk_{1}\max\left\{1,\frac{\alpha_{1}^{2}}{k_{1}}\right\}}=\frac{nk_{1}+1}{12}\frac{12}{nk_{1}+1}\sqrt{n{k_{1}}\max\left\{1,\frac{\alpha_{1}^{2}}{k_{1}}\right\}}.$
Now, by applying Lemma 2.3 with $m=nk_{1}+1$ and $r=12$, we get
(4.2) $\ \ \ \ \ \ \ \ \ \ \ \ \tan\theta_{c}\Big(nk_{1}+1,\alpha\Big)<\frac{12}{nk_{1}+1}\tan\theta_{c}\left(12,\,\frac{12}{nk_{1}+1}\sqrt{n{k_{1}}\max\left\{1,\frac{\alpha_{1}^{2}}{k_{1}}\right\}}\right)$
when $nk_{1}+1>12$. Note that, as $n\uparrow\infty$, the tangent term on the right hand side of (4.2) limits to $\tan 7.97^{\circ}$ according to the vanishing angle $\theta_{c}$ in the last column of the table on page 21 of [Law].
Therefore, when $n$ is large enough, by comparing leading term of (4.1) and (4.2) we have
$\theta_{c}\left(nk_{1}+1,\alpha\right)<\tan\theta_{c}\big(nk_{1}+1,\alpha\big)<\frac{R(L_{1}^{\times n})}{2}.$
According to Lawlor’s criterion Theorem 2.2, the cone $C(L_{1}^{\times n})$ is area-minimizing in $\mathbb{R}^{n(N_{1}+1)}$. $\square$
Remark 4.1.
To get full generality in Theorem 1.1, notice that finitely many $\{L_{i}\}$ also imply the finiteness property of $\max_{1\leq i\leq n}\left\{1,\frac{\alpha_{i}^{2}}{k_{i}}\right\}$ and consequently
(4.3) $\ \ \ \ \ \ \ \ \ \ \ \ \tan\theta_{c}\Big(k+1,\alpha\Big)<\frac{12}{k+1}\tan\theta_{c}\left(12,\,\frac{12}{k+1}\sqrt{k\max_{1\leq i\leq n}\left\{1,\frac{\alpha_{i}^{2}}{k_{i}}\right\}}\right)$
where $k$ is the dimension of a minimal product among $\{L_{i}\}$. Also note that (3.13) leads to $\cos R\leq 1-\frac{1}{k}\min_{1\leq i\leq n}\left\{1-\cos R_{i}\right\}$ which gives
(4.4) $R\geq\arccos\left(1-\frac{1}{k}\min_{1\leq i\leq n}\left\{1-\cos R_{i}\right\}\right)=\sqrt{2\min_{1\leq i\leq n}\left\{1-\cos R_{i}\right\}}\cdot\frac{1}{\sqrt{k}}+$...
As a result, whenever $k$ is sufficiently large one can apply Theorem 2.2 for the conclusion of Theorem 1.1.
If, further beyond finiteness, there are infinitely many $\{L_{\ell}\}_{\ell\in\mathscr{A}}$ with
(4.5) $\sup_{\ell\in\mathscr{A}}\left\{1,\frac{\alpha_{\ell}^{2}}{k_{\ell}}\right\}<\infty\text{ \ \ \ \ and \ \ \ \ }\inf_{\ell\in\mathscr{A}}\left\{k_{\ell}(1-\cos R_{\ell})\right\}>0,$
then the conclusion of Theorem 1.1 still holds.
Note that, by [L-Z], for fixed $k_{1}\geq 3$ and every $\delta,\delta^{\prime}>0$ there exists spiral minimal product type $L_{1}$ of dimension $k_{1}$ in some sphere with its normal radius $R_{1}<\delta$ and $\frac{\alpha_{1}^{2}}{k_{1}}>(\delta^{\prime})^{-1}$ where $\alpha_{1}$ is its supremum norm of second fundamental form with respect to unit normals. So generally there is no control no these two quantities for a minimal submanifold (of codimension larger than two) in sphere. However, machinery Theorem 1.1 works regardless of the initial figure of $L_{1}$ at all.
Apparently, the idea here can lead to many general configuration results. Among others, here we would like to mention a corollary of the above proof.
Corollary 4.2.
Let $L$ be a closed smooth minimal subamanifold in some $\mathbb{S}^{n}$. Then the cone over minimal product $L\times\mathbb{S}^{d}$ is area-minimizing for sufficiently large $d$.
Remark 4.3.
There exists integer $d_{0}$ such that the cone over the minimal product $L^{k}\times\mathbb{S}^{d_{1}}\times\cdots\mathbb{S}^{d_{\ell}}$ is area-minimizing whenever $d=\sum_{i=1}^{\ell}d_{i}\geq d_{0}$. If $L$ is a minimal isoparametric hypersurface or a focal submanifold, then this essentially falls into Theorem 4.18 in below and $k+d\geq 36$ would be sufficient for the area-minimality.
Remark 4.4.
Given a closed minimal hypersurface $L$ in $\mathbb{S}^{n}$ ($n\geq 3$) and any integer $\tau\geq 2$. By the corollary and remark, it follows that based on $L$ there exist infinitely many area-minimizing cones in Euclidean spaces having fixed codimension $\tau$.
Now we are ready to give a proof of Theorem 1.2.
Proof of Theorem 1.2. Set
[…]
and
[…]
Then the mapping
${\tt F}:B\in\mathscr{B}\longrightarrow B\cap\text{the corresponding unit sphere}\in\mathscr{A}$
is injective.
Conversely, for every $(\subset\mathbb{S}^{N_{A}}) 𝒜\in\mathscr{A}$, according to Corollary 4.2 there exists a smallest integer $d_{A}\geq\dim A+1$ such that the cone over $A\times\mathbb{S}^{d}$ is area-minimizing when $d\geq d_{A}$. So, if we define
(4.6) ${\tt G}:A\longrightarrow C\left(A\times\mathbb{S}^{d_{A}+N_{A}}\right),$
then it is not hard to see that $\tt G$ is injective from $\mathscr{A}$ to $\mathscr{B}$.
Hence, by the Schröder–Bernstein theorem in the set theory, there exists a bijective map between $\mathscr{A}$ and $\mathscr{B}$. As a result, they share the same cardinality. $\square$
Remark 4.5.
Here we add $N_{A}$ in the dimensional part of $\mathbb{S}$ in (4.6) to distinguish $A$ being a minimal submanifold in different ambient spheres, since we regard $A\subset\mathbb{S}^{N_{A}}$ and $A\subset\mathbb{S}^{N_{A}+1}$ as two distinct elements in $\mathscr{A}$.
Remark 4.6.
By [CM], for any positive integer $\tt n$, there is a real $\tt n$-dimensional family of minimal Lagrangian tori in $\mathbb{C}P^{2}$ and therefore an $\tt n$-dimensional family of special Legendrian tori in $\mathbb{S}^{5}$. Hence the cardinality of special Legendrian tori in $\mathbb{S}^{5}$ is no less than that of the real numbers.
In order to get more configuration results in the category of Lawlor cones, we dive into more details about the minimal product construction and the vanishing angle controls in the rest part of the paper.
4.2. Key control on normal radius of minimal product construction.
We wish to show that, in some circumstances, the minimal product of links $L_{1},L_{2}$ of two Type-c cones can again span a Type-c cone. So we aim to get the inequality $\theta_{c}(k+1,\alpha)<\frac{R(L_{1}\times L_{2})}{2}$ where $k=k_{1}+k_{2}$ and $k_{i}$ is the dimension of $L_{i}$ for $i=1,2$.
In this subsection we give some structural inequality for normal radius of minimal products. The first three results actually work for constant products, see Remark 3.3. Here the constant product construction merely means that $\{\lambda_{i}\}$ in the algorithm (3.1) are nonzero numbers satisfying $\sum_{i=1}^{n}\lambda_{i}^{2}=1$.
Lemma 4.7.
Let $R_{1},R_{2},R$ be normal radii of $L_{1},L_{2}$ and $L_{1}\times L_{2}$ respectively with $k_{1},k_{2}\geq 1$. Then $R\leq\frac{\pi}{2}$.
Proof.
Adding up the two possible values in (3.12), we have
(4.7) $2\cos R\geq 1+\lambda_{1}^{2}\cos R_{1}+\lambda_{2}^{2}\cos R_{2}\geq 0.$
Hence $R\leq\frac{\pi}{2}$ and equality holds exactly when $R_{1}=R_{2}=\pi$ and $k_{1}=k_{2}$ simultaneously. ∎
Lemma 4.8.
$\cos R>\cos R_{i}$ and $0<\sin R<\sin R_{i}$ for $i=1,2$.
Proof.
Since $\cos R=\max\{\lambda_{2}^{2}+\lambda_{1}^{2}\cos R_{1},\,\lambda_{1}^{2}+\lambda_{2}^{2}\cos R_{2}\}\geq\cos R_{i}$ for $i=1,2$. As function $\cos(\cdot)$ is strictly decreasing in the interval $0,\frac{\pi}{2}]$, by Lemma 4.7 we have $\sin R\leq\sin R_{i}$ for $i=1,2$. Since any closed submanifold has positive normal radius, we get $\sin R>0,\sin R_{i}>0$ and all inequalities in the proof must be strict ones. ∎
The following is an immediate corollary.
Corollary 4.9.
Assume that (3.12) is taken for $i=1$, i.e., $\cos R=1-\lambda_{1}^{2}(1-\cos R_{1})$. Then, we have
(4.8) $\tan\frac{R}{2}>\lambda_{1}^{2}\tan\frac{R_{1}}{2}.$
Proof.
It follows from Lemma 4.8 that
$\tan\frac{R}{2}=\frac{1-\cos R}{\sin R}>\frac{\lambda_{1}^{2}(1-\cos R_{1})}{\sin R_{1}}=\lambda_{1}^{2}\tan\frac{R_{1}}{2}.$
Here are some thoughts useful for our next step. Suppose that $L_{1}$ spans a Type-c cone. Then
$\tan\frac{R}{2}>\lambda_{1}^{2}\tan\frac{R_{1}}{2}>\lambda_{1}^{2}\tan\theta_{c}(k_{1}+1,\alpha_{1}),$
and we would be happy if, for some reason,
$\tan\theta_{c}(k_{1}+k_{2}+1,\alpha)\leq\frac{k_{1}}{k_{1}+k_{2}}\tan\theta_{c}(k_{1}+1,\alpha_{1}),$
since this would imply that $\theta_{c}(k+1,\alpha)<\frac{R(L_{1}\times L_{2})}{2}$ as we want. But, comparing this with Lemma 2.3, one can see that the ratio outside $\tan$ is $\frac{k_{1}+1}{k_{1}+k_{2}+1}$ which is larger than the constant $\frac{k_{1}}{k_{1}+k_{2}}$. It turns out that it could be very hard or likely impossible to improve the ratio on the right hand side of (2.4) in Lemma 2.3 to be $\frac{k_{1}}{k_{1}+k_{2}}$. (The same problem exists for the Type-F estimate (2.5) of Lemma 2.4.) To conquer this difficulty, instead we turn to improve the inequality (4.8) and get what we need for our purpose.
Lemma 4.10 (Key control).
As in Corollary 4.9, $\cos R=1-\lambda_{1}^{2}(1-\cos R_{1})$. Assume that (1) $k_{2}\geq 4$ or (2) $k_{2}\geq 3,k_{1}\geq 2$ and assume that $R_{1}\leq\frac{\pi}{2}$. Then for the minimal product construction we have
(4.9) $\tan\frac{R}{2}\geq\frac{k_{1}+1}{k_{1}+k_{2}+1}\tan\frac{R_{1}}{2}.$
Proof.
With $k=k_{1}+k_{2}$, the assumption tells that $\cos^{2}R=\left(\frac{k_{2}+k_{1}\cos R_{1}}{k}\right)^{2}$. Hence,
$k^{2}\sin^{2}R$ $=$ $k^{2}-k_{2}^{2}-2k_{1}k_{2}\cos R_{1}-k_{1}^{2}\cos^{2}R_{1}$
$=$ $k^{2}-k_{2}^{2}-k_{1}^{2}-2k_{1}k_{2}\cos R_{1}+k_{1}^{2}\sin^{2}R_{1}$
$=$ $2k_{1}k_{2}(1-\cos R_{1})+k_{1}^{2}\sin^{2}R_{1}$
$\leq$ $2k_{1}k_{2}(1-\cos R_{1})(1+\cos R_{1})+k_{1}^{2}\sin^{2}R_{1}$
$=$ $k_{1}(2k_{2}+k_{1})\sin^{2}R_{1}$
and thus $\sin R\leq\frac{\sqrt{k_{1}(2k_{2}+k_{1})}}{k}\sin R_{1}.$
Now, following the argument of Corollary 4.9, in order for
$\tan\frac{R}{2}\geq\frac{k_{1}}{\sqrt{k_{1}(2k_{2}+k_{1})}}\tan\frac{R_{1}}{2}\geq\frac{k_{1}+1}{k+1}\tan\frac{R_{1}}{2},$
it requires either $k_{2}\geq 4$ or $k_{2}\geq 3,k_{1}\geq 2$. ∎
Remark 4.11.
Similarly as in Remark 3.4, under the assumption that $\cos R(L_{1}\times\cdots\times L_{n})=1-\lambda_{1}^{2}(1-\cos R_{1})$ and $\sum_{i=2}^{n}k_{i}\geq 4$ and $R_{1}\leq\frac{\pi}{2}$, one can extend the result to the situation of multiple inputs
(4.10) $\tan\frac{R(L_{1}\times\cdots\times L_{n})}{2}\geq\frac{k_{1}+1}{\sum_{i=1}^{n}k_{i}+1}\tan\frac{R_{1}}{2}.$
Here we think of $L_{1}\times\cdots\times L_{n}$ as $L_{1}\times(L_{2}\times\cdots\times L_{n})$.
Remark 4.12.
Note that by Lemma 4.7 every minimal product type $L_{1}$ automatically satisfies the requirement $R_{1}\leq\frac{\pi}{2}$.
4.3. Configuration results in the category of Typc-c Lawlor cones.
For better exposition we introduce the following concepts.
Definition 4.13.
Let $\alpha$ be the supremum norm square of second fundamental form with respect to unit normals for closed minimal submanifold $L$ of dimension $k$ in sphere. We define the quantity $\frac{\alpha^{2}}{k}$ to be the slope of $L$, denoted by s(L)(L). Say $L$ (and also C(L)) is of class (a), (b), (c) if s(L)<1(L)<1, s(L)=1(L)=1 and s(L)>1(L)>1 respectively. The union of classes (a) and (b) is called class (I), and that for (b) and (c) is called class (II).
There are plentiful examples for each type. For some neat ones, we may consider isoparametric foliations of spheres. For brief information about their s(L)(L), readers are referred to [TZ]. Interested readers may check references therein for further knowledge.
A closed embedded hypersurface in a unit sphere is called isoparametric if it has constant principal curvatures. This property is inherited by its parallel hypersurfaces. In such way it decides an isoparametric foliation with two focal submanifolds $M_{+}$ and $M_{-}$. Let g be the number of distinct principal curvatures, $\cot\theta_{\beta}$ $(\beta=1,\cdots,g;0<\theta_{1}<\cdots<\theta_{\tt g}<\pi)$ of a regular isoparametric leaf and $m_{\beta}$ the multiplicity of $\cos\theta_{\beta}$. Münzner showed that g must be 1,2,3,4 or 6 and $m_{\beta}=m_{\beta+2}$ with indices mod g. Distributions of $({\tt g},m_{1},m_{2})$ are completely clear. The slopes for the unique minimal isoparametric hypersurface and two focal submanifolds in an isomparametric foliation of sphere only depend on the distribution $({\tt g},m_{1},m_{2})$ regardless of a complete classification of isoparametric foliations of spheres.
[…]
Here g=1 for totally geodesic hyperspheres with normal radius $\pi$, which belong to class (a); g=2 for Clifford type minimal hypersurfaces, which belong to class (b); g=3 has four examples with $m_{1}=m_{2}=m_{3}=1,2,4,8$, which belong to class (c); for g=6 there are only two values $m_{1}=m_{2}=1,2$ and corresponding hypersurfaces are of class (c). When g=4 there are infinitely many geometrically distinct examples of minimal hypersurfaces and all of them are of class (c).
Example 2. Focal submanifolds. For any isoparametric foliation with g=4, note that $M_{+}$ has dimension $m_{1}+2m_{2}$ and $M_{-}$ has dimension $2m_{1}+m_{2}$ with $\alpha_{+}^{2}=2m_{2}$ and $\alpha_{-}^{2}=2m_{1}$ respectively. Hence, both are of class (a).
[…]
Note that each focal submanifolds has normal radius exactly $\frac{2\pi}{\tt g}$.
Here are some simple properties of these classes, according to Lemma 3.1.
Proposition 4.14.
Minimal product $L_{1}\times L_{2}$ of two minimal submanifolds $L_{1},L_{2}$ in spheres belongs to either class of (b) or (c).
Proposition 4.15.
Given $L_{1},L_{2}$ of class (I). Then $L_{1}\times L_{2}$ belongs to class (b).
Proposition 4.16.
If either $L_{1}$ or $L_{2}$ is of class (c), then $L_{1}\times L_{2}$ belongs to class (c). If either $L_{1}$ or $L_{2}$ is of class (II), then $L_{1}\times L_{2}$ belongs to class (II).
Example 3. Lawlor classified all area-minimizing cones over links which are minimal products of spheres. All these links are of class (b).
Example 4. Our Theorem 1.1 can manufacture a plethora of elements of class (b) and class (c). For example, given finitely many closed minimal submanifolds $\{L_{i}\}$ in spheres, whenever they are not much curved, i.e., $\alpha_{i}^{2}\leq k_{i}$, our Theorem 1.1 can apply and produce lots of members in class (b).
Moreover, to the countably many $\{L_{\ell}\}_{\ell\in\mathscr{A}}$ of focal submanifolds with ${\tt g}=4$ in Example 2 above which satisfy (4.5), we can apply Remark 4.1.
In fact as an application we can cover the main result of [TZ] based on the structural discovery found here. For doing so let us recall the following basic property.
Proposition 4.17.
Let $L_{1}^{k_{1}}$ be either a minimal isoparametric hypersurface or a focal submanifold and $R_{1}$ its normal radius. Then it follows that
(4.11) $k_{1}(1-\cos R_{1})>0.8\text{ \ \ \ and \ \ \ }\alpha_{1}^{2}\leq 5k_{1}.$
Proof.
When $L_{1}$ is a minimal isoparametric hypersurface of a sphere with $\tt g$}, the value of $\cos R_{1}$ is clear and the author computed that in (56) of [TZ] :
(4.12) $\cos R_{1}=\begin{cases}-1&\ {\tt g}=1,\\ 1-\frac{2\min\{m_{1},\;m_{2}\}}{m_{1}+m_{2}}&\ {\tt g}=2,\\ \frac{1}{2}&\ {\tt g}=3,\\ \sqrt{1-\frac{\min\{m_{1},\;m_{2}\}}{m_{1}+m_{2}}}&\ {\tt g}=4,\\ \frac{\sqrt{3}}{2}&\ {\tt g}=6.\end{cases}$
Note that we have now $k_{1}=m_{1}+m_{2}$ for ${\tt g}=2$; $k_{1}=3m_{1}$ for ${\tt g}=3$; $k_{1}=2(m_{1}+m_{2})$ for ${\tt g}=4$; $k_{1}=6m_{1}$ for ${\tt g}=6$. So,
(4.13) $k_{1}(1-\cos R_{1})=\begin{cases}2k_{1}&\ {\tt g}=1,\\ 2\min\{m_{1},\;m_{2}\}&\ {\tt g}=2,\\ \frac{3}{2}m_{1}&\ {\tt g}=3,\\ 2(m_{1}+m_{2})\left(1-\sqrt{1-\frac{\min\{m_{1},\;m_{2}\}}{m_{1}+m_{2}}}\right)&\ {\tt g}=4,\\ 6m_{1}\left(1-\frac{\sqrt{3}}{2}\right)&\ {\tt g}=6.\end{cases}$
The second last is larger than $\min\{m_{1},\;m_{2}\}$ and the last is bigger than $0.8038m_{1}$.
When $L_{1}$ is a focal submanifold of some isoparametric foliation of sphere, as already mentioned $R_{1}=\frac{2\pi}{\tt g}$. Note that we have $k_{1}=0$ for ${\tt g}=1$ (not in our consideration with positive $k_{1}$); $k_{1}=m_{1}$ or $m_{2}$ for ${\tt g}=2$; $k_{1}=2m_{1}$ for ${\tt g}=3$; $k_{1}=2m_{1}+m_{2}$ or $m_{1}+2m_{2}$ for ${\tt g}=4$; $k_{1}=5m_{1}$ for ${\tt g}=6$. So, $k_{1}(1-\cos R_{1})\geq 1$.
As a result, the quantity $k_{1}(1-\cos R_{1})\geq 0.8$ as stated.
As for $\alpha_{1}^{2}$, it was also computed in [TZ] with the conclusion $\alpha_{1}^{2}=({\tt g}-1)k_{1}$ for the minimal isoparametric hypersurfaces; $\alpha_{1}^{2}=0$ for focal submanifolds with ${\tt g}=2$; $\alpha_{1}^{2}=\frac{k_{1}}{3}$ for focal submanifolds with ${\tt g}=3$; $\alpha_{1}^{2}=2m_{2}$ or $2m_{1}\leq k_{1}-1$ for focal submanifolds with ${\tt g}=4$; $\alpha_{1}^{2}=\frac{4k_{1}}{3}$ for focal submanifolds with ${\tt g}=6$. Hence, $\alpha_{1}^{2}\leq 5k_{1}$. ∎
Theorem 4.18.
Let $\{L_{\ell}\}_{\ell\in\mathscr{A}}$ in Remark 4.1 be the collection of all minimal isoparametric hypersurfaces and focal submanifolds of spheres. Then cones, of dimension no less than $37$, over minimal products among the collection are all area-minimizing.
Proof.
Use $L$ to denote the minimal product of dimension $k>11$ with normal radius $R$. As
$\tan\theta_{c}\Big(k+1,\alpha\Big)$ $\leq$ $\frac{12}{k+1}\tan\theta_{c}\left(12,\,\frac{12}{k+1}\sqrt{5k}\right)$
when $k\geq 36$ $<$ $\frac{12}{k+1}\tan\theta_{c}\left(12,\,\sqrt{19}\right)$
$<$ $\frac{12}{k+1}\tan 11.23^{\circ}$
$<$ $\frac{2.383}{k+1}$
and by Proposition 4.17
$R>\arccos\left(1-\frac{0.8}{k}\right)=\sqrt{\frac{1.6}{k}}+\text{ positive terms}>\sqrt{\frac{1.6}{k}}\,$,
now let us solve the sufficient inequality to apply Lawlor’s criterion
$\theta_{0}(C(L))<\theta_{c}(k+1,\alpha)<\tan\theta_{c}\Big(k+1,\alpha\Big)<\frac{2.383}{k+1}<\frac{\sqrt{1.6}}{2\sqrt{k}}<\frac{R}{2}\,.$
It suffices to have $14.2*k<(k+1)^{2}$ and this is always true for $k\geq 13$. Therefore, the theorem gets proved by the requirement (4.3). ∎
A useful lemma for further discussion is the following.
Lemma 4.19 (Existence of $\theta_{c}$).
Suppose $L_{1},L_{2}$ span Type-c cones with $k_{1}+k_{2}\geq 11$. Then, for $\alpha=\alpha(L_{1}\times L_{2})$, the vanishing angle $\theta_{c}(k_{1}+k_{2}+1,\alpha)$ exists.
Remark 4.20.
By no means that $11$ is optimal, we simply take the advantage of the vanishing angle table of [Law]. One can definitely improve the sum requirement. Since successive minimal products make dimension larger and larger, so any concrete lower bound on total dimension would be fine for configuration results in below.
Proof.
Note that by Lemma 2.3, when $\alpha^{2}=k=k_{1}+k_{2}$,
(4.15) $\tan\theta_{c}\Big(k+1,\sqrt{k}\Big)\leq\frac{12}{k+1}\tan\theta_{c}\left(12,\,\frac{12}{k+1}\sqrt{k}\right)<\frac{12}{k+1}\tan\theta_{c}\left(12,\sqrt{12}\right)$
exists and, when $\alpha^{2}=\frac{k\alpha_{\tau}^{2}}{k_{\tau}}$ either for $\tau=1$ or for $\tau=2$,
(4.16) $\tan\theta_{c}\Big(k+1,\alpha\Big)$
$<$ $\frac{k_{\tau}+1}{k+1}\tan\theta_{c}\left(k_{\tau}+1,\,\frac{k_{\tau}+1}{k+1}\alpha\right)$
$=$ $\frac{k_{\tau}+1}{k+1}\tan\theta_{c}\left(k_{\tau}+1,\frac{\sqrt{k_{\tau}}+\frac{1}{\sqrt{k_{\tau}}}}{\sqrt{k}+\frac{1}{\sqrt{k}}}\alpha_{\tau}\right)$
$<$ $\frac{k_{\tau}+1}{k+1}\tan\theta_{c}\left(k_{\tau}+1,\alpha_{\tau}\right)$
and the latter exists as by assumption either $C(L_{\tau})$ is a Type-c cone.
Thus, no matter which value $\alpha^{2}$ takes from $\left\{k,k\frac{\alpha^{2}_{1}}{k_{1}},k\frac{\alpha^{2}_{2}}{k_{2}}\right\}$, we see from the above that $\theta_{c}(k+1,\alpha)$ always exists. ∎
Our first configuration result in this subsection is the following.
Theorem 4.21.
Suppose $L_{1},L_{2}$ span Type-c cones with $k_{1},\,k_{2}\geq 3$. If $(L_{1})=s$$(L_{2})\geq 1$, then their minimal product $L_{1}\times L_{2}$ spans a Type-c cone of class (II).
Remark 4.22.
If $(L_{1})=s$$(L_{2})<1$, the proof in below will bother (4.15) with additional assumption $k_{1}+k_{2}\geq 11$ or some smaller number, see Remark 4.20.
Proof.
Without loss of generality, assume that $\cos R=1-\lambda_{1}^{2}(1-\cos R_{1})$. There are two cases to be discussed separately.
Case 1. $R_{1}\leq\frac{\pi}{2}$. Now we apply the key control Lemma 4.10 under the assumption on dimensions ($k_{1},\,k_{2}\geq 3$ would be enough for (4.3), as condition (2) of Lemma 4.10 satisfied) to get
$\tan\frac{R}{2}$ $\geq$ $\frac{k_{1}+1}{k_{1}+k_{2}+1}\tan\frac{R_{1}}{2}$
$\geq$ $\frac{k_{1}+1}{k_{1}+k_{2}+1}\tan\big(\theta_{c}(k_{1}+1,\alpha_{1})\big)$
$>$ $\tan\big(\theta_{c}(k_{1}+k_{2}+1,\frac{k_{1}+k_{2}+1}{k_{1}+1}\sqrt{{\tt s}(L_{1})k_{1}})\big)$
$=$ $\tan\big(\theta_{c}(k_{1}+k_{2}+1,\frac{\sqrt{k_{1}+k_{2}}+\frac{1}{\sqrt{k_{1}+k_{2}}}}{\sqrt{k_{1}}+\frac{1}{\sqrt{k_{1}}}}\sqrt{{\tt s}(L_{1})(k_{1}+k_{2})})\big)$
(4.18) $>$ $\tan\big(\theta_{c}(k_{1}+k_{2}+1,\alpha)\big)$
The last inequality (4.18) is due to the increasing property of $x+\frac{1}{x}$ when $x\geq 1$, the increasing property of $\theta_{c}(m,\cdot)$ in the second slot  [^3] This is clear in [Law] but to be self-contained we give a short proof in appendix., and the fact that $\alpha^{2}=(k_{1}+k_{2}){\tt s}(L_{1})$ under the assumption.
Case 2. $\frac{\pi}{2}<R_{1}\leq\pi$. By the proof of Lemma 4.19 for $(L_{1})=s$$(L_{2})\geq 1$ the vanishing angle $\theta_{c}(k+1,\alpha)$ always exists. Note that (4.15) is not bothered so that we have no assumption on total dimension. Moreover, there is a free rough control $\theta_{c}(k+1,\alpha)\leq\tan\theta_{c}(k+1,\alpha)<\frac{1}{\alpha}$ since the function $(1-\alpha t)e^{\alpha t}$ vanishes at $t=\frac{1}{\alpha}$ (see appendix for more information about the equivalent calibration inequality where solution needs to reach zero before $t=\frac{1}{\alpha}$).
Under the assumption of Case 2, we have $\cos R\leq 1-\lambda_{1}^{2}$ which implies
$2\left(\frac{R}{2}\right)^{2}\geq 2\sin^{2}{\frac{R}{2}}=1-\cos R\geq\lambda_{1}^{2},$
and hence $\frac{R}{2}\geq\sqrt{\frac{k_{1}}{2k}}$. As $\alpha\geq\sqrt{k}$ , we also get $\frac{R}{2}\geq\theta_{c}(k+1,\alpha)$ when $k_{1}\geq 2$.
Therefore, in both Case 1 and Case 2, one can apply Lawlor’s criterion and the proof gets completed. ∎
Remark 4.23.
Examples of class (b) are many, see Example 4. All elements of class (b) has slope one. This theorem shows its power and derives a universal structure for the category of Type-c cones of class (b).
Not as in Theorem 1.1 where sufficiently many copies of a given $L$ are needed, for links of Type-c cones of class (II) two copies can work.
Corollary 4.24.
Let $L^{k}$ be the link of a Type-c cone of class (II) with $k\geq 3$ and $n\geq 2$. Then the minimal product $L^{\times n}$ can also span a Type-c cone of class (II).
From the proof of Theorem 4.21, we know that as long as $\cos R$ and $\alpha^{2}$ are determined by the same slot of the minimal product then the minimal product spans a Type-c cone. There are many specific cases for which this coincidence in the same slot occurs. A general structure result is the following (cf. Corollary 4.2).
Theorem 4.25.
Let $L^{k}$ be the link of a Type-c cone of class (II) and $n\geq k\geq 3$. Then $L\times\mathbb{S}^{n}$ can span a Type-c cone.
Proof.
Under the dimension assumption, we can see from (3.12) that
$\cos R=1-\frac{k}{n+k}(1-\cos R(L)).$
Moreover, by Lemma 3.1 we have $\alpha^{2}(L\times\mathbb{S}^{n})=(k+n)\max\{1,{\tt s}(L)\}=(k+n){\tt s}(L)$. Consequently, following the reasoning in the preceding paragraph, the conclusion stands true. ∎
Moreover, a further structure result related to Theorem 1.1 can be obtained.
Theorem 4.26.
Suppose $L_{1}^{\times s_{1}}\times\cdots\times L_{n}^{\times s_{n}}$ with $n\geq 2$ and $s_{i}\geq 1$ for $i=1,\cdots,n$ spans a Type-c cone. Then so does $\tilde{L}\triangleq L_{1}^{\times(s_{1}+\ell_{1})}\times\cdots\times L_{n}^{\times(s_{n}+\ell_{n})}$ where $\ell_{i}\geq 0$ and $\sum_{i+1}^{n}k_{i}\ell_{i}\geq 3$.
Remark 4.27.
Here $L_{i}$ itself may not be area-minimizing.
Remark 4.28.
This result in particular indicates that, for given $\{L_{i}\}_{i=1}^{n}$, Theorem 1.1 can guarantee the existence of uniform $\tilde{k}$ such that whenever total dimension is no less than $\tilde{k}$ the cone C(L~)C($\tilde${L}) is then a Type-c cone and thus area-minimizing. When each $k_{i}\geq 3$, this $\tilde{k}$ could be less than $k$ in Remark 4.1 in general.
Proof.
Without loss of generality, we consider $\check{L}\triangleq L_{1}^{\times(s_{1}+1)}\times\cdots\times L_{n}^{\times s_{n}}$ with $k_{1}\geq 3$ only and other situations can be derived accordingly. Note that by Lemma 4.7 it follows $R(L)\leq\frac{\pi}{2}$ where $L\triangleq L_{1}^{\times s_{1}}\times\cdots\times L_{n}^{\times s_{n}}$. By Proposition 4.14 or Lemma 3.1 we have ${\tt s}(L)\geq\max\{1,{\tt s}(L_{1})\}$. Therefore, both $R(\check{L})$ and ${\tt s}(\check{L})$ are decided by the first input $L$ of the minimal product of $L$ and $L_{1}$. When $k_{1}\geq 3$, the key control Lemma 4.10 applies and the argument in the proof in Case 1 of Theorem 4.21 leads us to the conclusion of the theorem. ∎
In fact more subtle structures can be established. To illustrate the point, let us mention merely two more situations, based on which interested readers may figure out many more structures.
Theorem 4.29.
Let $L^{k}$ be the link of a Type-c cone with $k\geq 3$. If $L$ is of class (II) with no other assumption on dimension or of class (a) with $d=k+\sum_{i=1}^{n}d_{i}\geq 11$, then $L^{k}\times\mathbb{S}^{d_{1}}\times\cdots\mathbb{S}^{d_{n}}$ spans a Type-c cone.
Proof.
The reason is following. If some slot $\mathbb{S}^{d_{\tau}}$ decides $R$ instead of component $L$, then we are in Case 2 in the proof of Theorem 4.21 and can get $\frac{R}{2}\geq\sqrt{\frac{d_{\tau}}{d}}$ now and $\theta_{c}(d+1,\alpha)<\frac{1}{\alpha}\leq\sqrt{\frac{d_{\tau}}{d}}$ by the existence of $\theta_{c}$ when $d\geq 11$. If $L$ is of class (II), the requirement on total dimension can be dropped. ∎
Instead of successively multiplying spheres, one can also consider using other kinds of links of class (I) (or their mixed types) for universal configuration results. For example, let $\tilde{M}$ be a focal submanifolds with ${\tt g}=4$ of sufficiently large dimension $n$. Since normal radius of each of them is always $\frac{\pi}{2}$, we get the following.
Theorem 4.30.
Let $L^{k}$ be the link of a Type-c cone of class (II) and $\frac{n}{2}\geq k\geq 3$. Then $L\times\tilde{M}^{n}$ can span a Type-c cone.
Remark 4.31.
One can take $L$ to be a resulting Type (II) cone from Theorem 1.1. Then Theorems 4.25, 4.29 and 4.30 explore some extra structures. Moreover, if $L$ is of class (b), then so are the resultings in these theorems.
Remark 4.32.
We would like to point out that Lemma 4.19, Theorem 4.21, Remarks 4.23, Corollary 4.24, Theorems 4.25, 4.29 and 4.30 will also be true if “Type-c” is replaced by “Type-F”. The key point is to verify (4.16) and (4.18) with the extra factor from (2.5). We provide details in appendix.
Remark 4.33.
When “Type-c” is replaced by “Type-F”, the requirement the total dimension $\geq 11$ in Lemma 4.19, Remark 4.22 and Theorem 4.29 can be replaced by the total dimension $\geq 7$, since corresponding upper bound of replaced (4.15) exists as $\theta_{F}(8,\sqrt{7})<16^{\circ}$ according to Lawlor’s table in [Law]. Also see $(\ast)$ in appendix for remaining the requirement $k_{1},k_{2}\geq 3$” unchanged in the Type-F version of Theorem 4.21.
Configuration results of this paper can provide a great number of local models in [z0] for observing duality obstruction of calibrations in smooth category.
5. Appendix
We collect proofs of Lemma 2.4, Footnote 3 and Remark 4.32 in appendix.
5.1. Proof of Lemma 2.4.
As briefly mentioned in §2 that Lawlor’s construction of area-decreasing projection is somehow equivalent to an inequality for the dual form to be a calibration. With $t=\tan\theta$ and $r$ the dimension of the cone, the inequality for calibration form is the following
$\left(g-\frac{tg^{\prime}}{r}\right)^{2}+\left(\frac{g^{\prime}}{r}\right)^{2}\leq\det\left(\textbf{I}-t\,\textbf{h}_{ij}^{v}\right)^{2},\text{ \ \ \ \ \ \ with }g(0)=1\ \ \text{ and \ \ }g\text{\ reaches }0.$
As Lawlor uses a uniform curve $\gamma$, the above inequality becomes
$\left(g-\frac{tg^{\prime}}{r}\right)^{2}+\left(\frac{g^{\prime}}{r}\right)^{2}\leq\inf_{x\in L,\,v\in S_{x}}\det\left(\textbf{I}-t\,\textbf{h}_{ij}^{v}\right)^{2}$
and being a Type-F cone means that there exists $g(t)$ such that
(5.1) $\left(g-\frac{tg^{\prime}}{r}\right)^{2}+\left(\frac{g^{\prime}}{r}\right)^{2}\leq\left(F(\alpha,t,r)\right)^{2},\text{ \ \ \ \ \ \ with }g(0)=1\ \ \text{ and \ \ }g\text{\ reaches }0.$
Recalling (2.3), we want to compare
(5.2) $F(\alpha,t,r)=\left(1-\alpha t\sqrt{\frac{r-1}{r}}\right)\left(1+\frac{\alpha t}{\sqrt{r(r-1)}}\right)^{r-1},$
and
(5.3) $F(\beta,s,m)=\left(1-\beta s\sqrt{\frac{m-1}{m}}\right)\left(1+\frac{\beta s}{\sqrt{m(m-1)}}\right)^{m-1}.$
Set $x=\alpha t\sqrt{\frac{r-1}{r}}=\beta s\sqrt{\frac{m-1}{m}}$ and they have simpler expressions
(5.4) $\left(1-x\right)\left(1+\frac{x}{{r-1}}\right)^{r-1}\text{\ \ \ and \ \ \ }\left(1-x\right)\left(1+\frac{x}{{m-1}}\right)^{m-1}.$
Note that, with fixed $x>0$,
(5.5) $\frac{d}{dn}\left(\log\left(1+\frac{x}{n}\right)^{n}\right)=\log\left(1+\frac{x}{n}\right)+n\cdot\frac{-\frac{x}{n^{2}}}{1+\frac{x}{n}}=\log\left(1+\frac{x}{n}\right)-\frac{x}{n+x}.$
Now taking derivative with respect to $x$ we have
$\frac{d}{dx}\left\{\log\left(1+\frac{x}{n}\right)-\frac{x}{n+x}\right\}=\frac{\frac{1}{n}}{1+\frac{x}{n}}+\frac{-n-x+x}{(n+x)^{2}}=\frac{1}{n+x}-\frac{n}{(n+x)^{2}}=\frac{x}{n+x}>0.$
As (5.5) is zero at $x=0$, we know (5.5) is positive when $x>0$. Hence the latter in (5.4) is larger than the former since $m>r$.
Inspired by the proof of Lemma 2.3 in [Law], we choose $\beta$ such that
$t=\frac{\beta}{\alpha}s\sqrt{\frac{(m-1)r}{m(r-1)}}=\frac{m}{r}s,$
namely,
$\beta\sqrt{\frac{(m-1)r}{m(r-1)}}=\frac{m}{r}{\alpha}.$
Set $h(s)=g(t)=g(\frac{m}{r}s)$. Then $\dot{h}=\frac{m}{r}g^{\prime}(\frac{m}{r}s)$. It follows from (5.1) for $g,r$ and $\alpha$ that
$\left(h(s)-\frac{s\dot{h}}{r}\right)^{2}+\left(\frac{\dot{h}}{m}\right)^{2}\leq\left(F(\alpha,t,r)\right)^{2}\leq\left(F(\beta,s,m)\right)^{2}.$
Since $\dot{h}\leq 0\leq h$, we have
$\left(h(s)-\frac{s\dot{h}}{m}\right)^{2}+\left(\frac{\dot{h}}{m}\right)^{2}\leq\left(F(\beta,s,m)\right)^{2}.$
This means that $h$ solves the calibration inequality and hence, by $t=\frac{m}{r}s$, we get
(5.6) $\tan\left(\theta_{F}\left(m,\frac{m}{r}\sqrt{\frac{(r-1)m}{(m-1)r}}\alpha\right)\right)<\frac{r}{m}\tan\left(\theta_{F}\left(r,\alpha\right)\right).$
To compare with (2.4) one can move the extra factor (inverse) to the right hand side and end up with Lemma 2.4.
5.2. Proof of Footnote 3.
To show Footnote 3 about the monotonicity property for $\theta_{c}(r,\cdot)$ in the second slot, one just needs to observe that $(1-\alpha t)e^{\alpha t}$ in (2.2) is decreasing in $\alpha$. So, suppose that that
$\left(h(s)-\frac{s\dot{h}}{r}\right)^{2}+\left(\frac{\dot{h}}{r}\right)^{2}\leq\left((1-\beta s)e^{\beta s}\right)^{2}$
holds. By setting $t=s$ and $g(t)=h(s)$, when $\beta>\alpha$ one has
$\left(g(t)-\frac{t\dot{g}}{r}\right)^{2}+\left(\frac{g^{\prime}}{r}\right)^{2}\leq\left((1-\beta s)e^{\beta s}\right)^{2}<\left((1-\alpha t)e^{\alpha t}\right)^{2}.$
Therefore, $\theta_{F}(r,\alpha)<\theta_{F}(r,\beta)$.
Remark 5.1.
The increasing monotonicity for $\theta_{F}(r,\cdot)$ follows similary.
5.3. Proof of Remark 4.32.
Note that to have (4.18) for Type-F we need to check
$\sqrt{\frac{(k_{1}+k_{2}+1)k_{1}}{(k_{1}+k_{2})(k_{1}+1)}}\frac{\sqrt{k_{1}+k_{2}}+1\big/\sqrt{k_{1}+k_{2}}}{\sqrt{k_{1}}+1\big/\sqrt{k_{1}}}>1$
which is equivalent to
$\sqrt{\frac{k_{1}+k_{2}+1}{k_{1}+1}}\frac{1+\frac{1}{k_{1}+k_{2}}}{1+\frac{1}{k_{1}}}>1.$
For this we check whether $\frac{(x+1)^{3}}{x^{2}}$ is increasing in $x$. By differentiating $3\log(x+1)-2\log x$, we have $\frac{3}{x+1}-\frac{2}{x}=\frac{x-2}{x(x+1)}>0$ when $x>2$. Hence the remark gets proved $(\ast)$ when $k_{1}\geq 2$ or when $k_{1}=1$ and $k_{2}\geq 4$.
Due to exactly the same reason (4.16) and with extra factor added also holds.