### v5

# Some configuration results for area-minimizing cones

Abstract.
We discover some very general configuration results for constructing area-minimizing cones. In particular, given any closed minimal submanifold in some Euclidean sphere, every cone over the minimal product of sufficiently many copies of the submanifold turns out to be area-minimizing; meanwhile every cone over the minimal product of the submanifold and a round sphere of sufficiently large dimension is also area-minimizing. Here no additional geometric assumption (e.g. on isometry group or second fundamental form) is required. Moreover, we establish that the category of regular area-minimizing cones in Euclidean spaces and that of closed minimal submanifolds in Euclidean spheres share the same cardinality.
Key words and phrases:
Regular area-minimizing cone, Lawlor criterion
1. Introduction
Let $L$ be a set in the unit sphere $\mathbb{S}^{N}$ in Euclidean space $(\mathbb{R}^{{N+1}},g_{E})$ and $C(L)$ be the cone $\{tx:t\geq 0,\ x\in L\}$ over its link $L$. It is called area-minimizing in geometric measure theory, if the truncated cone $C\bigcap\mathbb{B}^{{N+1}}(1)$ has least mass among all integral currents (see [FF]) with boundary $L$. A fundamental result (Theorem 5.4.3 in [F]) asserts the existence of tangent cone at every point of an area-minimizing rectifiable current and moreover each tangent cone itself area-minimizing. Although ambient Riemannian manifold can be quite arbitrary, after blowing-up those cones always live in standard Euclidean spaces. Lots of efforts have been made in studying important properties and finding examples of them in various situations, e.g., [JS, BdGG, BL, HL, HS, FK, Ch, Law0, Law, NS, XYZ2, z, TZ] and etc.
When $L$ is an embedded closed smooth submanifold, the cone $C(L)$ is smooth away from the origin and we say it is a regular cone. In this paper, we search for regular area-minimizing cones and end up with very general configuration results.
Let us get started with a basic question regarding the structure of the category of regular area-minimzing cones.
$(\star)$ Given two regular area-minimizing cones $C(L_{1})$ and $C(L_{2})$ in $\mathbb{R}^{{N_{1}+1}}$ and $\mathbb{R}^{{N_{2}+1}}$ respectively, is the cone over minimal product $L_{1}\times L_{2}$ still area-minimizing in $\mathbb{R}^{{N_{1}+N_{2}+2}}$?
Here the minimal product of $L_{1}$ and $L_{2}$ of dimension $k_{1}$ and $k_{2}$ respectively is
(1.1) $\left(\sqrt{\frac{k_{1}}{{k_{1}+k_{2}}}}L_{1},\sqrt{\frac{k_{2}}{{k_{1}+k_{2}}}}L_{2}\right)\subset\mathbb{S}^{N_{1}+N_{2}+1}.$
Apparently, the minimal product can apply to multiple inputs inductively. Along this line, Lawlor (Theorem 5.1.1, [Law]) completely settled down $(\star)$ when each input $L_{i}$ is $\mathbb{S}^{N_{i}}$, with $i=1,2,\cdots,n$ and $N_{i}\in\mathbb{Z}_{+}$. In particular, whenever $\Sigma_{i=1}^{n}N_{i}>7$, the corresponding cone is area-minimizing. In recent work [TZ], cones of dimension no less than 37 over minimal products of the category of typical minimal submanifolds associated to isoparametric foliations of spheres are confirmed area-minimizing. From these, one can observe that the dimension plays an important role for $(\star)$ and that some minimal submanifolds in spheres themselves cannot span area-minimizing cones but their minimal products do.
In this paper, among others we get the following.
Theorem 1.1.
Given embedded closed minimal submanifolds $\{L_{i}\}$ respectively in Euclidean spheres with $i=1,2,\cdots,n$, every cone over the minimal product $L$ of sufficiently many copies among these $\{L_{i}\}$ is area-minimizing.
Remark 1.2.
Every closed minimal submanifold with no other assumptions at all can be used as a brick to construct area-minimizing cones in different combinatorial ways. The only payment is that the resulting area-minimizing cone may have high dimension.
Another neat result is the following.
Theorem 1.3.
Let $L$ be a closed smooth minimal subamanifold in some $\mathbb{S}^{N}$. Then the cone over minimal product $L\times\mathbb{S}^{d}$ is area-minimizing for sufficiently large $d$.
Remark 1.4.
In fact there exists integer $d_{0}$ such that the cone over the minimal product $L^{k}\times\mathbb{S}^{d_{1}}\times\cdots\mathbb{S}^{d_{\ell}}$ is area-minimizing whenever $d=\sum_{i=1}^{\ell}d_{i}\geq d_{0}$. If $L$ is a minimal isoparametric hypersurface or a focal submanifold, then this falls into [TZ] and $k+d\geq 36$ would be sufficient for the area-minimality.
Prior to this paper, all known regular area-minimizing cones are tightly related to either group actions or the theory of isoparametric foliations on spheres. The above theorems in particular break these restrictions. Moreover, a corollary of the theorems from the point of set theory is established as follow.
Corollary 1.5.
Regular area-minimizing cones in Euclidean spaces and closed minimal submanifolds in Euclidean spheres share the same cardinality.
The main tool is Lawlor’s curvature criterion for a regular cone to be area-minimizing. To be self-contained we briefly explain his method in §2. Some useful information about minimal product will be given in §3. Finally, the proofs of the theorems and corollary and more subtle results partially answering $(\star)$ and complementing the theorems are collected in §4.
Acknowledgments. The author wishes to thank Professor Blaine Lawson for drawing his attention to the powerful work [Law] of Professor Gary Lawlor. This work was sponsored in part by NSFC (Grant Nos. 12022109, 11971352).
2. Lawlor’s curvature criterion
Area-minimizing hypercone is a key ingredient to solve the famous Bernstein problem [B] about entire minimal graph of codimension one, see [fle, de2, A]. It was Jim Simons who found first non-trivial stable minimal hypercones $C(S^{k}\times S^{k})$ with $k\geq 3$ in [JS] and suspected that they are area-minimizing. In the following year, breakthrough was made in [BdGG] which confirmed those Simons cones’ area-minimality and based on this discovery counterexamples of non-planar minimal graphs to the Bernstein problem were found over domain $\mathbb{R}^{N}$ when $N\geq 8$.
In [HS] a beautiful characterization of area-minimizing hypercones was established. In each side of an area-minimizing hypercone, there is a unique dilation-invariant foliation by minimal hypersurfaces. Therefore, the unit norm vector fields lead to a foliation by their integral curves on $\mathbb{R}^{N+1}\sim 0$, which is also invariant under dilation. It naturally induces a dual co-degree one calibration form (see [HL]) and an area-nonincreasing projection (for surfaces of codimension one) to the hypercone in $\mathbb{R}^{N+1}\sim 0$.
For a given regular minimal cone, how to determine whether it is area-minimizing or not? The most effective way is to employ the theory of calibration. However, in general, it is not easy to directly find a desired calibration with mild singularity. Lawlor successfully developed a sufficient criterion (sometimes also necessary) looking for area-nonincreasing projection structure in suitable angular neighborhood of the minimal cone instead of in the whole space $\mathbb{R}^{N}$. If the boundary of the neighborhood is mapped to the origin under the projection, then one can send everything outside the neighborhood to the origin. In this way an area-nonincreasing projection can be globally defined.
To understand Lawlor’s criterion, let us recall the following with $k<N$.
Definition 2.1.
Let $L^{k}\subset\mathbb{S}^{N}$ be a closed embedded submanifold, $x\in L$ and $S_{x}$ denote the set of all unit normals of $L$ at $x$. Then the normal radius of $L$ at $x$ is
$R_{x}\triangleq\sup\left\{\,t\in\mathbb{R}_{+}\,:\,\exp_{x}(sv_{x})\not\in L\,\text{ for all }v_{x}\in S_{x}\text{ and }s\in(0,t)\right\}$
where $\exp_{x}(sv_{x})=(\cos s)\,x+(\sin s)\,v_{x}$ forms a normal geodesic along $v_{x}$. The quantity $R(L)\triangleq\inf\{\,R_{x}\,:\,x\in L\,\}$ is called normal radius of $L$.
For gaining area-minimizing projection, Lawlor considered in each normal wedge
$W_{x}\left(\frac{R_{x}}{2}\right)=C\left(\exp_{x}\left(\left[0,\frac{R_{x}}{2}\right]\cdot S_{x}\right)\right)$
[…]
(2.1) $\frac{dr}{d\theta}\leq r\sqrt{r^{2k+2}\cos^{2k}\theta\inf_{x\in L,\,v\in S_{x}}\left(\det\left(\textbf{I}-(\tan\theta)\textbf{h}_{ij}^{v}\right)\right)^{2}-1}$
with $r(0)=\|x\|=1$. Here $\textbf{h}_{ij}^{v}$ means the second fundamental form at $x$ with respect to $v$. To get narrowest curve $\gamma_{x}$, one needs to require that equality holds everywhere in (2.1) and wants to have the smallest $\theta_{0}$ such that $\lim_{\theta\uparrow\theta_{0}}r(\theta)=\infty$. As a result, the boundary of the normal wedge $W_{x}\left(\theta_{0}\right)$ is projected to the origin and global area-decreasing projection gets set up. Such $\theta_{0}$, if existed, is called vanishing angle of $C(L)$. Note that there is no focal point of $L$ in the neighborhood $\bigcup_{x\in L}W_{x}\left(\theta_{0}\right)$ since otherwise that focal point leads to a strictly negative expression under the square root symbol in (2.1).
Lawlor’s curvature criterion is the following.
Theorem 2.2 (Theorem 1.2.1, [Law]).
Given a regular minimal cone $C(L)\subset\mathbb{R}^{N+1}$, if the vanishing angle $\theta_{0}$ exists and $\theta_{0}\leq\frac{R(L)}{2}$, then $C(L)$ is area-minimizing.
Where does curvature play a role? How to check whether a given cone satisfies $\theta_{0}\leq\frac{R(L)}{2}$? In practice, Lawlor employed
(2.2) $(1-\alpha t)e^{\alpha t}<F(\alpha,t,k+1)\leq\det\left(\textbf{I}-t\,\textbf{h}_{ij}^{v}\right)$
where $t=\tan\theta$, $\alpha=\sup_{x\in L,v\in S_{x}}\|\textbf{h}_{ij}^{v}\|$ is the curvature term and
(2.3) $F(\alpha,t,k+1)=\left(1-\alpha t\sqrt{\frac{k}{k+1}}\right)\left(1+\frac{\alpha t}{\sqrt{k(k+1)}}\right)^{k},$
to replace the infimum term in (2.1) deriving two kinds of vanishing angles
$\theta_{c}(k+1,\alpha)>\theta_{F}(k+1,\alpha)\geq\theta_{0}(C(L)).$
So, if either $\theta_{F}$ or $\theta_{c}$ exists of finite value, then consequently vanishing angle $\theta_{0}$ must exist. For details readers are referred to [Law].
An important property about $\theta_{c}$ is the following, which provides an effective control on $\theta_{c}$.
Lemma 2.3 (Proposition 1.4.2, [Law]).
For $m,r\in\mathbb{Z}_{+}$ and $m>r$, it follows that
(2.4) $\tan\left(\theta_{c}\Big(m,\frac{m}{r}\alpha\Big)\right)<\frac{r}{m}\tan\Big(\theta_{c}(r,\alpha)\Big)\,.$
In fact a ratio type control on $\theta_{F}$ can also be established.
Lemma 2.4.
For $m,r\in\mathbb{Z}_{+}$ and $m>r$, it follows that
(2.5) $\tan\left(\theta_{F}\Big(m,\frac{m}{r}\alpha\Big)\right)<\frac{r}{m}\tan\left(\theta_{F}\left(r,\alpha\sqrt{\frac{(m-1)r}{(r-1)m}}\right)\right)\,.$
Although $F$ and $\theta_{F}$ are more accurate and efficient for applying the criterion, however in certain sense Lemma 2.4 is not as good as Lemma 2.3 since the extra factor $\sqrt{\frac{(m-1)r}{(r-1)m}}$ is larger than one. There will be no much difference in the scope of current paper when $m$ is sufficiently large and we save its proof in appendix.
Definition 2.5.
We call a regular minimal cone Type-c (or Type-F) Lawlor cone if it can be verified area-minimizing using the vanishing angle $\theta_{c}$ (or $\theta_{F}$), i.e., $\theta_{c}\leq\frac{R(L)}{2}$ (or $\theta_{F}\leq\frac{R(L)}{2}$).
We will show that the situations in Theorems 1.1 and 1.3 bring us Type-c cones. Besides, some considerations regarding $(\star)$ will be made in §4 (and results work for Type-F as well).
3. Minimal product
The (constant) minimal product  [^2] We say constant minimal product to distinguish this from spiral minimal product in [L-Z]. in §1 can naturally induce a nice algorithm to generate minimal submanifold for multiple minimal submanifolds
(3.1) $L_{1}\times\cdots\times L_{n}\triangleq\left(\lambda_{1}L_{1},\cdots,\lambda_{n}L_{n}\right)\subset\mathbb{S}^{N_{1}+\cdots+N_{n}+n-1}.$
where $\lambda_{i}=\sqrt{\frac{k_{i}}{k}}$ with $k_{i}$ the dimension of $L_{i}$ and $k=\sum_{i=1}^{n}k_{i}>0$. In this paper we shall focus on the case that every $k_{i}>0$.
There have been many specific situations where minimal product has been employed. For all $L_{i}$ being full spheres, the structure produces Clifford type minimal surfaces. When these $L_{i}$ are R-spaces, [OS] studied whether cones over minimal products of them are area-minimizing. For general minimal submanifolds as inputs, see [X] for a proof through the Takahashi Theorem. Around 2016 it was rediscovered independently by [TZ] and [CH] using different methods. Reader may choose either familiar way to understand the structure.
3.1. Formula of $\alpha^{2}$ for minimal product.
For simplicity, we mainly focus on $n=2$ in the paper and all related results for $n\geq 3$ can be gained by induction. In [TZ], information about the second fundamental form of minimal product has been clearly figured out. Let $k_{1},k_{2}$ be the dimension of $L_{1},L_{2}$ respectively. Take $x_{1}\in L_{1}$ and $x_{2}\in L_{2}$. Together with $\eta_{0}=(\lambda_{2}x_{1},-\lambda_{1}x_{2})$, orthonormal bases $\{\sigma_{1},\cdots,\sigma_{N_{1}-k_{1}}\}$ and $\{\tau_{1},\cdots,\tau_{N_{2}-k_{2}}\}$ of $T_{x_{1}}^{\perp}L_{1}$ and $T_{x_{2}}^{\perp}L_{2}$ lead to an orthonormal basis
$\{(\sigma_{1},0),\cdots,(\sigma_{N_{1}-k_{1}},0),\;(0,\tau_{1}),\cdots,(0,\tau_{N_{2}-k_{2}}),\;\eta_{0}\}$
of the normal space of $L_{1}\times L_{2}$ at point $(\lambda_{1}x_{1},\lambda_{2}x_{2})$ in $\mathbb{S}^{N_{1}+N_{2}+1}$.
Let $A$ stand for the symbol of shape operators. Then with respect to natural tangential orthonormal bases, we have the following relations:
[…]
Let $\alpha_{1}^{2},\,\alpha_{2}^{2},\,\alpha^{2}$ be the supremum norm squares of corresponding second fundamental forms with respect to unit normals for $L_{1},\ L_{2}$ and $L_{1}\times L_{2}$ respectively. Then the following is obvius.
Lemma 3.1.
As marked in the above, it follows that, for $k_{1},k_{2}>0$,
$\alpha^{2}=(k_{1}+k_{2})\max\left\{\frac{\alpha_{1}^{2}}{k_{1}},\,\frac{\alpha_{2}^{2}}{k_{2}},\,1\right\}.$
Proof.
Note that the matrices of shape operators with respect to three subclasses of unit normals $\{(\sigma_{i},0)\}$, $\{(0,\tau_{i})\}$ and $\{\eta_{0}\}$ are mutually perpendicular. ∎
3.2. Formula of normal radius for minimal product
Another vital quantity in Lawlor’s criterion is the normal radius, so in this subsection we derive its formula rigorously for the minimal product structure.
Let $R_{1},\,R_{2},\,R$ be the normal radii of $L_{1},\ L_{2}$ and $L_{1}\times L_{2}$ in corresponding Euclidean spheres respectively. For $(\lambda_{1}x_{1},\lambda_{2}x_{2})$ in $L_{1}\times L_{2}$, the combination
(3.2) $\nu=a_{1}(\xi_{1},0)+a_{2}(0,\xi_{2})+a_{0}\eta_{0}$
where $\xi_{1}$, $\xi_{2}$ are unit normals of $L_{1}$ and $L_{2}$ at $x_{1}$ and $x_{2}$ respectively, and $a_{1}^{2}+a_{2}^{2}+a_{0}^{2}=1$, can exhaust all unit normals of $L_{1}\times L_{2}$ at $(\lambda_{1}x_{1},\lambda_{2}x_{2})$. Then, in $\mathbb{S}^{N_{1}+N_{2}+1}$, each unit normal induces a normal geodesic
$\exp_{(\lambda_{1}x_{1},\lambda_{2}x_{2})}(s\nu)$
$=$ $(\cos s)(\lambda_{1}x_{1},\lambda_{2}x_{2})+(\sin s)\nu$
$=$ $\big((\lambda_{1}\cos s+a_{0}\lambda_{2}\sin s)x_{1}+(a_{1}\sin s)\xi_{1},$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ (\lambda_{2}\cos s-a_{0}\lambda_{1}\sin s)x_{2}+(a_{2}\sin s)\xi_{2}\big)$
Therefore, $R$ is the smallest positive number such that there exists
$p\triangleq(\lambda_{1}x_{1},\lambda_{2}x_{2})\in L_{1}\times L_{2}$
satisfying $q\triangleq(\lambda_{1}y_{1},\lambda_{2}y_{2})=\exp_{p}(R\nu_{p})\in L_{1}\times L_{2}$ for some $\nu_{p}\in S_{p}$. Thus,
(3.4) $(\lambda_{1}\cos R+a_{0}\lambda_{2}\sin R)x_{1}+(a_{1}\sin R)\xi_{1}=\lambda_{1}\big((\cos\phi_{1})x_{1}+(\sin\phi_{1})\xi_{1}\big)$
(3.5) $(\lambda_{2}\cos R-a_{0}\lambda_{1}\sin R)x_{2}+(a_{2}\sin R)\xi_{2}=\lambda_{2}\big((\cos\phi_{2})x_{2}+(\sin\phi_{2})\xi_{2}\big)$
for some $\phi_{1},\phi_{2}\in\mathbb{R}$ such that
$y_{1}=(\cos\phi_{1})x_{1}+(\sin\phi_{1})\xi_{1}\text{\ \ \ and \ \ \ }y_{2}=(\cos\phi_{2})x_{2}+(\sin\phi_{2})\xi_{2}.$
Somehow similarly as argued in [TZ], there are three possibilities to realize normal radius $R$, but notice that, a priori, possible values of $\phi_{1}$ and $\phi_{2}$ in the above may not be discrete as in the isoparametric situation in [TZ]. Moreover, except being a point of sphere, the normal radius of a closed submanifold in sphere is no larger than $\pi$.
The formula of normal radius for minimal product can be derived.
Lemma 3.2.
Let $R_{1}$ and $R_{2}$ be the normal radii of $L_{1}$ and $L_{2}$ in $\mathbb{S}^{N_{1}}$ and $\mathbb{S}^{N_{2}}$ respectively. Then the normal radius $R\in(0,\pi)$ of $L_{1}\times L_{2}$ in $\mathbb{S}^{N_{1}+N_{2}+1}$ satisfies
(3.6) $\cos R=1-\min\left\{\lambda_{1}^{2}(1-\cos R_{1}),\,\lambda_{2}^{2}(1-\cos R_{2})\right\}.$
Remark 3.3.
Note that the result is in fact valid when $\lambda_{1}$ and $\lambda_{2}$ are just two nonzero real numbers $a_{1},a_{2}$ with $a_{1}^{2}+a_{2}^{2}=1$. However, we will restrict ourselves in the minimal product construction unless otherwise specified.
Remark 3.4.
If some $L_{i}$ is the full sphere $\mathbb{S}^{N_{i}}$, then $R_{i}$ is defined to be $\pi$.
Proof.
The three possibilities for $R$ based on (3.4) and (3.5) are the followings.
Case I. $y_{1}\neq x_{1}$ and $y_{2}\neq x_{2}$.
[…]
(3.7) $\cos R=\lambda_{1}^{2}\cos\phi_{1}+\lambda_{2}^{2}\cos\phi_{2}\leq\lambda_{1}^{2}\cos R_{1}+\lambda_{2}^{2}\cos R_{2}$
Case II. $y_{1}\neq x_{1}$ and $y_{2}=x_{2}$ (or the other way around).
Now $\cos\phi_{2}=1$ and (3.7) becomes
(3.8) $\cos R\leq\lambda_{1}^{2}\cos R_{1}+\lambda_{2}^{2}.$
Case III. $y_{1}=x_{1}$ and $y_{2}=x_{2}$.
Since the ambient is Euclidean sphere, if $R$ occurs in this case then $R=2\pi$ (here $2\pi$ corresponds to normal geodesic curve for direction $\nu=\eta_{0}$ when neither $L_{1}$ nor $L_{2}$ contains antipodal points).
So, to search for the value of $R$ one should focus on Case II to see whether
(3.9) $\cos R=\lambda_{1}^{2}\cos R_{1}+\lambda_{2}^{2}.$
can be reached. Note that it is always true that
$0<R<\min\{\lambda_{1}R_{1},\,\lambda_{2}R_{2}\}<\pi.$
Hence $\sin R\neq 0$. So, in Case II, $a_{2}$ has to be zero in (3.5). Now we want to solve:
Given $\phi_{1}=R_{1}$ for some $x_{1}$ and direction $\xi_{1}$ realizing normal radius $R_{1}$ of $L_{1}$ in $\mathbb{S}^{N_{1}}$, can we get $a_{0},\,a_{1}$ (with $a_{0}^{2}+a_{1}^{2}=1$) and $\nu=a_{1}(\xi_{1},0)+a_{0}\eta_{0}$ to make (3.4), (3.5) and (3.9) hold?
This can be done by
(3.10) $a_{0}\sin R=\lambda_{1}\lambda_{2}(\cos R_{1}-1),$
$a_{1}\sin R=\lambda_{1}\sin R_{1}.\ \ \ \ \ \ \ \ \ \ \ \ \$
automatically with $a_{0}^{2}+a_{1}^{2}=1$.
Therefore $\cos R$ takes the larger value of $\lambda_{1}^{2}\cos R_{1}+\lambda_{2}^{2}$ and $\lambda_{1}^{2}+\lambda_{2}^{2}\cos R_{2}$ and (3.6) follows for genuine submanifold inputs.
Regarding Remark 3.4, when $L_{1}$ (or $L_{2}$ or both) is a full sphere, the normal radius $R_{1}$ (or $R_{2}$ or both) cannot be well defined as in Definition 2.1. For our purpose, we always define the normal radius of a full sphere $\mathbb{S}^{N_{1}}$ in itself to be $\pi$. The reason is the following. It can be realized as a totally geodesic sphere in $\mathbb{S}^{N_{1}+\ell}$ for $\ell>0$ and now $\mathbb{S}^{N_{1}}$ has normal radius $\pi$. So the above argument works for the minimal product and the last $\ell$ components of $\mathbb{R}^{N_{1}+1+\ell}$ can be completely dropped as they contribute nothing (let $\xi$ be a unit normal vector to the minimal product at some point and if $\xi$ has a nonzero part coming from the last $\ell$ components of $\mathbb{R}^{N_{1}+1+\ell}$ then the geodesic with initial velocity vector $\xi$ in $\mathbb{S}^{N_{1}+\ell+N_{2}}$ will run at least length $\pi$ to return to the minimal product and hence this will not realize the normal radius as the normal radius is strictly less than $\pi$ for the minimal product - not a totally geodesic sphere, actually no more than $\frac{\pi}{2}$ by Lemma 4.6).  [^3] For example, the normal radius of minimal product $\mathbb{S}^{N_{1}}\times\mathbb{S}^{N_{2}}$ in $\mathbb{S}^{N_{1}+N_{2}+1}$ is realized by $\eta_{0}$ and can reach $\frac{\pi}{2}$ only when $N_{1}=N_{2}$ according to (3.9), cf. the beautiful uniform formula in [Law] for normal radius of minimal product among full spheres. Thus, (3.6) still holds with the understanding in Remark 3.4. ∎
Remark 3.5.
Applying this to $L_{1}\times L_{2}\times L_{3}=(L_{1}\times L_{2})\times L_{3}$, we have
$\cos R(L_{1}\times L_{2}\times L_{3})$
$=$ $1-\min\left\{\frac{k_{1}+k_{2}}{k_{1}+k_{2}+k_{3}}(1-\cos R(L_{1}\times L_{2})),\,\frac{k_{3}}{k_{1}+k_{2}+k_{3}}(1-\cos R_{3})\right\}$
$=$ $1-\min\Bigg\{\frac{k_{1}}{k}(1-\cos R_{1}),\,\frac{k_{2}}{k}(1-\cos R_{2}),\,\frac{k_{3}}{k}(1-\cos R_{3})\Bigg\}.$
By induction, one can get the following formula.
Corollary 3.6.
Let $R_{i}$ be the normal radii of $L_{i}$ in $\mathbb{S}^{N_{i}}$ for $i=1,\cdots,n$. Then the normal radius $R(L_{1}\times\cdots\times L_{n})\in(0,\pi)$ of $L_{1}\times\cdots\times L_{n}$ in $\mathbb{S}^{N_{1}+\cdots+N_{n}+n-1}$ satisfies
(3.11) $\cos R(L_{1}\times\cdots\times L_{n})=1-\min_{1\leq i\leq n}\left\{\lambda_{i}^{2}(1-\cos R_{i})\right\}.$
Remark 3.7.
Similar statement formulated independently and some partial discussion included in [JCX].
4. Configuration results
This section consists of three parts.
In §4.1 the general configuration results Theorems 1.1 and 1.3 and Corollary 1.5 will be proved. The main idea is to gain $\theta_{c}<\frac{R}{2}$ in the asymptotic sense based on Lemmas 2.3 and 3.2.
In §4.2 some useful controls especially a key inequality on normal radius of minimal product $L_{1}\times L_{2}$ will be given in Lemma 4.9. Such an inequality together with Lemma 2.3 can lead to $\theta_{c}<\frac{R}{2}$ if both $\theta_{c}$ and $R$ of $L_{1}\times L_{2}$ are decided by the same component, which itself spans a Type-c cone. As a result, $L_{1}\times L_{2}$ also spans a Type-c cone.
In §4.3 we shall provide a partial positive answer to Question $(\star)$ in §1 under some assumptions. Besides, a quick proof of the main result of [TZ] will be explained in Theorem 4.17 following Remark 4.1. Moreover, we derive some more subtle structures (e.g. dimension gap of Theorem 1.1 in Theorem 4.23) complementing the asymptotic results in Theorems 1.1 and 1.3. Some other configuration results within the category of Type-c cones will also be discussed.
4.1. Asymptotic results for general closed minimal submanifold.
For simplicity, suppose that the set $\{L_{i}\}$ of closed minimal submanifolds in spheres in assumption contains exactly one element $L_{1}^{k_{1}}\subset\mathbb{S}^{N_{1}}$.
Proof of Theorem 1.1. Let $R_{1}>0$ be the normal radius of $L_{1}$ in $\mathbb{S}^{N_{1}}$. Denote the minimal product of $n$-copies of $L_{1}$ by $L_{1}^{\times n}$. Then, by (3.11) and Lemma 4.6 in below, the normal radius $R(L_{1}^{\times n})$ is no more than $\frac{\pi}{2}$ and its exact value is
(4.1) $\arccos\left(1-\frac{1}{n}(1-\cos R_{1})\right)=\sqrt{2(1-\cos R_{1})}\cdot\frac{1}{\sqrt{n}}+\text{ higher order terms}.$
On the other hand, denote by $\alpha_{1}$ and $\alpha$ the supremum norms of second fundamental forms of $L_{1}$ and $L_{1}^{\times n}$ respectively. By induction, Lemma 3.1 implies that
$\alpha=\sqrt{nk_{1}\max\left\{1,\frac{\alpha_{1}^{2}}{k_{1}}\right\}}=\frac{nk_{1}+1}{12}\frac{12}{nk_{1}+1}\sqrt{n{k_{1}}\max\left\{1,\frac{\alpha_{1}^{2}}{k_{1}}\right\}}.$
Now, by applying Lemma 2.3 with $m=nk_{1}+1$ and $r=12$, we get
(4.2) $\ \ \ \ \ \ \ \ \ \ \ \ \tan\theta_{c}\Big(nk_{1}+1,\alpha\Big)<\frac{12}{nk_{1}+1}\tan\theta_{c}\left(12,\,\frac{12}{nk_{1}+1}\sqrt{n{k_{1}}\max\left\{1,\frac{\alpha_{1}^{2}}{k_{1}}\right\}}\right)$
when $nk_{1}+1>12$. Note that, as $n\uparrow\infty$, the tangent term on the right hand side of (4.1) limits to $\tan 7.97^{\circ}$ according to the vanishing angle $\theta_{c}$ in the last column of the table on page 21 of [Law].
Therefore, when $n$ is large enough, by comparing leading terms of (4.1) and (4.1) we have
$\theta_{c}\left(nk_{1}+1,\alpha\right)<\tan\theta_{c}\big(nk_{1}+1,\alpha\big)<\frac{R(L_{1}^{\times n})}{2}.$
According to Lawlor’s criterion Theorem 2.2, the cone $C(L_{1}^{\times n})$ is area-minimizing in $\mathbb{R}^{n(N_{1}+1)}$. $\square$
Remark 4.1.
To get full generality in Theorem 1.1, notice that finitely many $\{L_{i}\}$ also imply the finiteness property of $\max_{1\leq i\leq n}\left\{1,\frac{\alpha_{i}^{2}}{k_{i}}\right\}$ and consequently
(4.3) $\ \ \ \ \ \ \ \ \ \ \ \ \tan\theta_{c}\Big(k+1,\alpha\Big)<\frac{12}{k+1}\tan\theta_{c}\left(12,\,\frac{12}{k+1}\sqrt{k\max_{1\leq i\leq n}\left\{1,\frac{\alpha_{i}^{2}}{k_{i}}\right\}}\right)$
where $k$ is the dimension of a minimal product among $\{L_{i}\}$. Also note that (3.11) leads to $\cos R\leq 1-\frac{1}{k}\min_{1\leq i\leq n}\left\{1-\cos R_{i}\right\}$ which gives
(4.4) $R\geq\arccos\left(1-\frac{1}{k}\min_{1\leq i\leq n}\left\{1-\cos R_{i}\right\}\right)=\sqrt{2\min_{1\leq i\leq n}\left\{1-\cos R_{i}\right\}}\cdot\frac{1}{\sqrt{k}}+$...
As a result, whenever $k$ is sufficiently large one can apply Theorem 2.2 for the conclusion of Theorem 1.1.
If, further beyond finiteness, there are infinitely many $\{L_{\ell}\}_{\ell\in\mathscr{A}}$ with
(4.5) $\sup_{\ell\in\mathscr{A}}\left\{1,\frac{\alpha_{\ell}^{2}}{k_{\ell}}\right\}<\infty\text{ \ \ \ \ and \ \ \ \ }\inf_{\ell\in\mathscr{A}}\left\{k_{\ell}(1-\cos R_{\ell})\right\}>0,$
then the conclusion of Theorem 1.1 still holds.
Note that, by [L-Z], for fixed $k_{1}\geq 3$ and every $\delta,\delta^{\prime}>0$ there exists spiral minimal product type $L_{1}$ of dimension $k_{1}$ in some sphere with its normal radius $R_{1}<\delta$ and $\frac{\alpha_{1}^{2}}{k_{1}}>(\delta^{\prime})^{-1}$ where $\alpha_{1}$ is its supremum norm of second fundamental form with respect to unit normals. So generally there is no control no these two quantities for a minimal submanifold (of codimension larger than two) in sphere. However, machinery works regardless of the initial figure of $L_{1}$ at all for spanning area-minimizing cones $C(L_{1}^{\times n})$ in the asymptotic sense in Theorem 1.1.
Apparently, the idea here can lead to many general configuration results. Among others, we get a proof of Theorem 1.3.
Proof of Theorem 1.3. Let $k_{1}$ be the dimension of $L\subset\mathbb{S}^{N}$. Then by Lemma 3.2 the normal radius $R=R(L\times\mathbb{S}^{d})$ of $L\times\mathbb{S}^{d}$ satisfies
$\cos R=1-\min\left\{\frac{k_{1}}{k_{1}+d}(1-\cos R_{1}),\,\frac{2d}{k_{1}+d}\right\}.$
Hence, when $d$ is large (e.g. $d\geq k_{1}$), we get
$\cos R=1-\frac{k_{1}}{k_{1}+d}(1-\cos R_{1})$
Similar to the argument in the proof of Theorem 1.1 we have
$\arccos\left(1-\frac{k_{1}}{k_{1}+d}(1-\cos R_{1})\right)=\frac{\sqrt{2k_{1}(1-\cos R_{1})}}{\sqrt{k_{1}+d}}+\text{ higher order terms}.$
Now note that by Lemma 3.1
$\alpha^{2}=(k_{1}+d)\max\left\{\frac{\alpha_{1}^{2}}{k_{1}},\,1\right\}.$
Therefore, according to Lemma 2.3 with $m=k_{1}+d+1$ and $r=12$, we get
$\tan\theta_{c}\Big(k_{1}+d+1,\alpha\Big)<\frac{12}{k_{1}+d+1}\tan\theta_{c}\left(12,\,\frac{12}{k_{1}+d+1}\sqrt{(k_{1}+d)\max\left\{1,\frac{\alpha_{1}^{2}}{k_{1}}\right\}}\right)$
when $k_{1}+d+1>12$.
So, when $d$ is large enough, by comparing leading terms we have
$\theta_{c}\left(k_{1}+d+1,\alpha\right)<\tan\theta_{c}\big(k_{1}+d+1,\alpha\big)<\frac{R(L\times\mathbb{S}^{d})}{2}.$
By Lawlor’s criterion Theorem 2.2 the cone $C(L\times\mathbb{S}^{d})$ is area-minimizing in $\mathbb{R}^{n+d+2}$.
If one replaces $\mathbb{S}^{d}$ by $\mathbb{S}^{d_{1}}\times\cdots\mathbb{S}^{d_{\ell}}$, then the generalization in Remark 1.4 holds due to the same kind of reasoning. $\square$
Remark 4.2.
Given a closed minimal hypersurface $L$ in $\mathbb{S}^{n}$ ($n\geq 3$) and any integer $\tau\geq 2$. By Theorem 1.3 and Remark 1.4, it follows that based on $L$ there exist infinitely many area-minimizing cones in Euclidean spaces having the fixed codimension $\tau$.
Beside the advantage of control on the precise codimension as in Remark 1.4, Theorem 1.3 can lead to a simple proof of Corollary 1.5.
Proof of Corollary 1.5. Set
[…]
and
[…]
Then the mapping
${\tt F}:B\in\mathscr{B}\longrightarrow B\cap\text{the corresponding unit sphere}\in\mathscr{A}$
is obviously injective.
Conversely, for every $(\subset\mathbb{S}^{N_{A}}) 𝒜\in\mathscr{A}$, according to Theorem 1.3 there exists a smallest integer $d_{A}\geq\dim A+1$ such that the cone over $A\times\mathbb{S}^{d}$ is area-minimizing when $d\geq d_{A}$  [^4] Also see Theorem 4.23 for a possible gap behavior.. So, if we define
(4.6) ${\tt G}:A\longrightarrow C\left(A\times\mathbb{S}^{d_{A}+N_{A}}\right),$
then it is not hard to see that $\tt G$ is injective from $\mathscr{A}$ to $\mathscr{B}$.
Hence, by the Schröder–Bernstein theorem in the set theory, there exists a bijective map between $\mathscr{A}$ and $\mathscr{B}$. As a result, they share the same cardinality. $\square$
Remark 4.3.
Here we add $N_{A}$ in the dimensional part of $\mathbb{S}$ in (4.6) to distinguish $A$ being a minimal submanifold in different ambient spheres, since we regard $A\subset\mathbb{S}^{N_{A}}$ and $A(\subset\mathbb{S}^{N_{A}})\subset\mathbb{S}^{N_{A}+1}$ as two distinct elements in $\mathscr{A}$.
Remark 4.4.
By [CM], for any positive integer $\tt n$, there is a real $\tt n$-dimensional family of minimal Lagrangian tori in $\mathbb{C}P^{2}$ and therefore an $\tt n$-dimensional family of special Legendrian tori in $\mathbb{S}^{5}$. Hence the cardinality of special Legendrian tori in $\mathbb{S}^{5}$ is no less than that of the real numbers.
Remark 4.5.
In fact one can also apply Theorem 1.1 for Corollary 1.5 but needs to be more careful in choosing corresponding $\tt G$ map.
Before closing this subsection, we would like to remark that all previously discovered area-minimizing cones C(L~)C($\tilde${L}) through Lawlor’s curvature criterion share the property (or restriction) that the curvature function $\alpha$ remains constant in the entire $\tilde{L}$, whereas in our Theorems 1.1 and 1.3 $\alpha$ can be a non-constant function over $\tilde{L}$.
4.2. Key control on normal radius of minimal product construction.
Instead of asymptotic behaviors in §4.1, we shall focus on some control regarding Question $(\star)$ of §1 in this subsection. More precisely, we wish to show that, in some circumstances, the minimal product of links $L_{1},L_{2}$ of two Type-c cones can again span a Type-c cone. So we aim to get the inequality $\theta_{c}(k+1,\alpha)<\frac{R(L_{1}\times L_{2})}{2}$ where $k=k_{1}+k_{2}$ and $k_{i}$ is the dimension of $L_{i}$ for $i=1,2$.
In particular we will give some structural inequality for normal radius of minimal products. The first three results actually work for constant products, see Remark 3.3. Here the constant product construction merely means that $\{\lambda_{i}\}$ in the algorithm (3.1) are nonzero numbers satisfying $\sum_{i=1}^{n}\lambda_{i}^{2}=1$.
Lemma 4.6.
Let $R_{1},R_{2},R$ be normal radii of $L_{1},L_{2}$ and $L_{1}\times L_{2}$ respectively with $k_{1},k_{2}\geq 1$. Then $R\leq\frac{\pi}{2}$.
Proof.
Adding up the two possible values in (3.6), we have
(4.7) $2\cos R\geq 1+\lambda_{1}^{2}\cos R_{1}+\lambda_{2}^{2}\cos R_{2}\geq 0.$
Hence $R\leq\frac{\pi}{2}$ and equality holds exactly when $R_{1}=R_{2}=\pi$ and $k_{1}=k_{2}$ simultaneously. ∎
Lemma 4.7.
$\cos R>\cos R_{i}$ and $0<\sin R<\sin R_{i}$ for $i=1,2$.
Proof.
Since $\cos R=\max\{\lambda_{2}^{2}+\lambda_{1}^{2}\cos R_{1},\,\lambda_{1}^{2}+\lambda_{2}^{2}\cos R_{2}\}\geq\cos R_{i}$ for $i=1,2$. As function $\cos(\cdot)$ is strictly decreasing in the interval $0,\frac{\pi}{2}]$, by Lemma 4.6 we have $\sin R\leq\sin R_{i}$ for $i=1,2$. Since any closed submanifold has positive normal radius, we get $\sin R>0,\sin R_{i}>0$ and all inequalities in the proof must be strict ones. ∎
The following is an immediate corollary.
Corollary 4.8.
Assume that (3.6) is taken for $i=1$, i.e., $\cos R=1-\lambda_{1}^{2}(1-\cos R_{1})$. Then, we have
(4.8) $\tan\frac{R}{2}>\lambda_{1}^{2}\tan\frac{R_{1}}{2}.$
Proof.
It follows from Lemma 4.7 that
$\tan\frac{R}{2}=\frac{1-\cos R}{\sin R}>\frac{\lambda_{1}^{2}(1-\cos R_{1})}{\sin R_{1}}=\lambda_{1}^{2}\tan\frac{R_{1}}{2}.$
Here are some thoughts useful for our next step. Suppose that $L_{1}$ spans a Type-c cone. Then
$\tan\frac{R}{2}>\lambda_{1}^{2}\tan\frac{R_{1}}{2}>\lambda_{1}^{2}\tan\theta_{c}(k_{1}+1,\alpha_{1}),$
and we would be happy if, for some reason,
$\tan\theta_{c}(k_{1}+k_{2}+1,\alpha)\leq\frac{k_{1}}{k_{1}+k_{2}}\tan\theta_{c}(k_{1}+1,\alpha_{1}),$
since this could imply that $\theta_{c}(k+1,\alpha)<\frac{R(L_{1}\times L_{2})}{2}$ as we want. But, comparing this with Lemma 2.3, one can see that the ratio outside $\tan$ is $\frac{k_{1}+1}{k_{1}+k_{2}+1}$ which is larger than the constant $\frac{k_{1}}{k_{1}+k_{2}}$. It seems very hard or likely impossible to improve the ratio on the right hand side of (2.4) in Lemma 2.3 to be $\frac{k_{1}}{k_{1}+k_{2}}$. (The same problem exists for the Type-F estimate (2.5) of Lemma 2.4.) To conquer this difficulty, instead we turn to improve the inequality (4.8) and get what we need for our purpose.
Lemma 4.9 (Key control).
As in Corollary 4.8, $\cos R=1-\lambda_{1}^{2}(1-\cos R_{1})$. Assume
(1) $k_{2}\geq 4$ or (2) $k_{2}\geq 3,k_{1}\geq 2$
and assume $R_{1}\leq\frac{\pi}{2}$. Then for the minimal product construction we have
(4.9) $\tan\frac{R}{2}\geq\frac{k_{1}+1}{k_{1}+k_{2}+1}\tan\frac{R_{1}}{2}.$
Proof.
With $k=k_{1}+k_{2}$, the assumption tells that $\cos^{2}R=\left(\frac{k_{2}+k_{1}\cos R_{1}}{k}\right)^{2}$. Hence,
$k^{2}\sin^{2}R$ $=$ $k^{2}-k_{2}^{2}-2k_{1}k_{2}\cos R_{1}-k_{1}^{2}\cos^{2}R_{1}$
$=$ $k^{2}-k_{2}^{2}-k_{1}^{2}-2k_{1}k_{2}\cos R_{1}+k_{1}^{2}\sin^{2}R_{1}$
$=$ $2k_{1}k_{2}(1-\cos R_{1})+k_{1}^{2}\sin^{2}R_{1}$
$\left(\text{by assumption }0<R_{1}\leq\frac{\pi}{2}\right)\ \ \ \ \ \ \displaystyle\leq$ $2k_{1}k_{2}(1-\cos R_{1})(1+\cos R_{1})+k_{1}^{2}\sin^{2}R_{1}$
$=$ $k_{1}(2k_{2}+k_{1})\sin^{2}R_{1}$
and thus $\sin R\leq\frac{\sqrt{k_{1}(2k_{2}+k_{1})}}{k}\sin R_{1}.$
Following the argument of Corollary 4.8 but now with $\sin R\leq\frac{\sqrt{k_{1}(2k_{2}+k_{1})}}{k}\sin R_{1}$, in order for
$\tan\frac{R}{2}\geq\frac{k_{1}}{\sqrt{k_{1}(2k_{2}+k_{1})}}\tan\frac{R_{1}}{2}\geq\frac{k_{1}+1}{k+1}\tan\frac{R_{1}}{2},$
it requires either $k_{2}\geq 4$ or $k_{2}\geq 3,k_{1}\geq 2$. ∎
Remark 4.10.
Similarly as in Remark 3.5, under the assumption that $\cos R(L_{1}\times\cdots\times L_{n})=1-\lambda_{1}^{2}(1-\cos R_{1})$ and $\sum_{i=2}^{n}k_{i}\geq 4$ and $R_{1}\leq\frac{\pi}{2}$, one can extend the result to the situation of multiple inputs
(4.10) $\tan\frac{R(L_{1}\times\cdots\times L_{n})}{2}\geq\frac{k_{1}+1}{\sum_{i=1}^{n}k_{i}+1}\tan\frac{R_{1}}{2}.$
Here we think of $L_{1}\times\cdots\times L_{n}$ as $L_{1}\times(L_{2}\times\cdots\times L_{n})$.
Remark 4.11.
Note that by Lemma 4.6 every minimal product type $L_{1}$ automatically satisfies the requirement $R_{1}\leq\frac{\pi}{2}$.
With the preparation of this subsection, we are ready to derive more configuration results in the next subsection.
4.3. More configuration results.
For better exposition, let us introduce some useful concepts.
Definition 4.12.
Let $\alpha$ be the supremum norm square of second fundamental form with respect to unit normals for closed minimal submanifold $L$ of dimension $k$ in sphere. We define the quantity $\frac{\alpha^{2}}{k}$ to be the slope of $L$, denoted by s(L)(L). Say $L$ (and also C(L)) is of class (a), (b), (c) if s(L)<1(L)<1, s(L)=1(L)=1 and s(L)>1(L)>1 respectively. The union of classes (a) and (b) is called class (I), and that for (b) and (c) is called class (II).
There are plentiful examples for each type. For some neat ones, we may consider isoparametric foliations of spheres. Here we mention some brief information about their s(L)(L). Interested readers may check [TZ] and references therein for further knowledge.
A closed embedded hypersurface in a unit sphere is called isoparametric if it has constant principal curvatures. This property is inherited by its parallel hypersurfaces. In such way it decides an isoparametric foliation with two focal submanifolds $M_{+}$ and $M_{-}$. Let g be the number of distinct principal curvatures, $\cot\theta_{\beta}$ $(\beta=1,\cdots,g;0<\theta_{1}<\cdots<\theta_{\tt g}<\pi)$ of a regular isoparametric leaf and $m_{\beta}$ the multiplicity of $\cos\theta_{\beta}$. Münzner showed that g must be 1,2,3,4 or 6 and $m_{\beta}=m_{\beta+2}$ with indices mod g. Distributions of $({\tt g},m_{1},m_{2})$ are completely clear. The slopes for the unique minimal isoparametric hypersurface and two focal submanifolds in an isomparametric foliation of sphere only depend on the distribution $({\tt g},m_{1},m_{2})$ regardless of a complete classification of isoparametric foliations of spheres.
[…]
Here g=1 for totally geodesic hyperspheres with normal radius $\pi$, which belong to class (a); g=2 for Clifford type minimal hypersurfaces, which belong to class (b); g=3 has four examples with $m_{1}=m_{2}=m_{3}=1,2,4,8$, which belong to class (c); for g=6 there are only two values $m_{1}=m_{2}=1,2$ and corresponding hypersurfaces are of class (c). When g=4 there are infinitely many geometrically distinct examples of minimal hypersurfaces and all of them are of class (c).
Example 2. Focal submanifolds. For any isoparametric foliation with g=4, note that $M_{+}$ has dimension $m_{1}+2m_{2}$ and $M_{-}$ has dimension $2m_{1}+m_{2}$ with $\alpha_{+}^{2}=2m_{2}$ and $\alpha_{-}^{2}=2m_{1}$ respectively. Hence, both are of class (a).
[…]
Note that each focal submanifolds has normal radius exactly $\frac{2\pi}{\tt g}$.
Here are some simple properties of these classes, according to Lemma 3.1.
Proposition 4.13.
Minimal product $L_{1}\times L_{2}$ of two minimal submanifolds $L_{1},L_{2}$ in spheres belongs to either class of (b) or (c).
Proposition 4.14.
Given $L_{1},L_{2}$ of class (I). Then $L_{1}\times L_{2}$ belongs to class (b).
Proposition 4.15.
If either $L_{1}$ or $L_{2}$ is of class (c), then $L_{1}\times L_{2}$ belongs to class (c). If either $L_{1}$ or $L_{2}$ is of class (II), then $L_{1}\times L_{2}$ belongs to class (II).
Example 3. Lawlor classified all area-minimizing cones over links which are minimal products of spheres. All these links are of class (b).
Example 4. Our Theorem 1.1 can manufacture a constellation of elements of class (b) and class (c). For example, given finitely many closed minimal submanifolds $\{L_{i}\}$ in spheres, whenever they are not much curved, i.e., $\alpha_{i}^{2}\leq k_{i}$, our Theorem 1.1 can apply and produce lots of members in class (b).
Moreover, to the countably many $\{L_{\ell}\}_{\ell\in\mathscr{A}}$ of focal submanifolds with ${\tt g}=4$ in Example 2 above which satisfy (4.5), we can apply Remark 4.1.
In fact as an application we can cover the main result of [TZ] based on the structural discovery found here. For doing so let us recall the following basic property.
Proposition 4.16.
Let $L_{1}^{k_{1}}$ be either a minimal isoparametric hypersurface or a focal submanifold and $R_{1}$ its normal radius. Then it follows that
(4.11) $k_{1}(1-\cos R_{1})>0.8\text{ \ \ \ and \ \ \ }\alpha_{1}^{2}\leq 5k_{1}.$
Proof.
When $L_{1}$ is a minimal isoparametric hypersurface of a sphere with $\tt g$}, the value of $\cos R_{1}$ is clear and the author computed that in (56) of [TZ] :
(4.12) $\cos R_{1}=\begin{cases}-1&\ {\tt g}=1,\\ 1-\frac{2\min\{m_{1},\;m_{2}\}}{m_{1}+m_{2}}&\ {\tt g}=2,\\ \frac{1}{2}&\ {\tt g}=3,\\ \sqrt{1-\frac{\min\{m_{1},\;m_{2}\}}{m_{1}+m_{2}}}&\ {\tt g}=4,\\ \frac{\sqrt{3}}{2}&\ {\tt g}=6.\end{cases}$
Note that we have now $k_{1}=m_{1}+m_{2}$ for ${\tt g}=2$; $k_{1}=3m_{1}$ for ${\tt g}=3$; $k_{1}=2(m_{1}+m_{2})$ for ${\tt g}=4$; $k_{1}=6m_{1}$ for ${\tt g}=6$. So,
(4.13) $k_{1}(1-\cos R_{1})=\begin{cases}2k_{1}&\ {\tt g}=1,\\ 2\min\{m_{1},\;m_{2}\}&\ {\tt g}=2,\\ \frac{3}{2}m_{1}&\ {\tt g}=3,\\ 2(m_{1}+m_{2})\left(1-\sqrt{1-\frac{\min\{m_{1},\;m_{2}\}}{m_{1}+m_{2}}}\right)&\ {\tt g}=4,\\ 6m_{1}\left(1-\frac{\sqrt{3}}{2}\right)&\ {\tt g}=6.\end{cases}$
The second last is larger than $\min\{m_{1},\;m_{2}\}$ and the last is bigger than $0.8038m_{1}$.
When $L_{1}$ is a focal submanifold of some isoparametric foliation of sphere, as already mentioned $R_{1}=\frac{2\pi}{\tt g}$. Note that we have $k_{1}=0$ for ${\tt g}=1$ (not in our consideration with positive $k_{1}$); $k_{1}=m_{1}$ or $m_{2}$ for ${\tt g}=2$; $k_{1}=2m_{1}$ for ${\tt g}=3$; $k_{1}=2m_{1}+m_{2}$ or $m_{1}+2m_{2}$ for ${\tt g}=4$; $k_{1}=5m_{1}$ for ${\tt g}=6$. So, $k_{1}(1-\cos R_{1})\geq 1$.
As a result, the quantity $k_{1}(1-\cos R_{1})\geq 0.8$ as stated.
As for $\alpha_{1}^{2}$, it was also computed in [TZ] with the conclusion $\alpha_{1}^{2}=({\tt g}-1)k_{1}$ for the minimal isoparametric hypersurfaces; $\alpha_{1}^{2}=0$ for focal submanifolds with ${\tt g}=2$; $\alpha_{1}^{2}=\frac{k_{1}}{3}$ for focal submanifolds with ${\tt g}=3$; $\alpha_{1}^{2}=2m_{2}$ or $2m_{1}\leq k_{1}-1$ for focal submanifolds with ${\tt g}=4$; $\alpha_{1}^{2}=\frac{4k_{1}}{3}$ for focal submanifolds with ${\tt g}=6$. Hence, $\alpha_{1}^{2}\leq 5k_{1}$. ∎
Theorem 4.17.
Let $\{L_{\ell}\}_{\ell\in\mathscr{A}}$ in Remark 4.1 be the collection of all minimal isoparametric hypersurfaces and focal submanifolds of spheres. Then cones, of dimension no less than $37$, over minimal products among the collection are all area-minimizing.
Proof.
Use $L$ to denote the minimal product of dimension $k>11$ with normal radius $R$. As
$\tan\theta_{c}\Big(k+1,\alpha\Big)$ $\leq$ $\frac{12}{k+1}\tan\theta_{c}\left(12,\,\frac{12}{k+1}\sqrt{5k}\right)$
when $k\geq 36$ $<$ $\frac{12}{k+1}\tan\theta_{c}\left(12,\,\sqrt{19}\right)$
$<$ $\frac{12}{k+1}\tan 11.23^{\circ}$
$<$ $\frac{2.383}{k+1}$
and by Proposition 4.16
$R>\arccos\left(1-\frac{0.8}{k}\right)=\sqrt{\frac{1.6}{k}}+\text{ positive terms}>\sqrt{\frac{1.6}{k}}\,$,
now let us solve the sufficient inequality to apply Lawlor’s criterion
$\theta_{0}(C(L))<\theta_{c}(k+1,\alpha)<\tan\theta_{c}\Big(k+1,\alpha\Big)<\frac{2.383}{k+1}<\frac{\sqrt{1.6}}{2\sqrt{k}}<\frac{R}{2}\,.$
It suffices to have $14.2*k<(k+1)^{2}$ and this is always true for $k\geq 13$. Therefore, the theorem gets proved by the requirement (4.3). ∎
Now let us prove a useful lemma on the existence of vanishing angle.
Lemma 4.18 (Existence of $\theta_{c}$).
Suppose $L_{1},L_{2}$ span Type-c cones with $k_{1}+k_{2}\geq 11$. Then, for $\alpha=\alpha(L_{1}\times L_{2})$, the vanishing angle $\theta_{c}(k_{1}+k_{2}+1,\alpha)$ exists.
Proof.
When $\alpha^{2}=k=k_{1}+k_{2}$, by Lemma 2.3 one can see that
(4.15) $\tan\theta_{c}\Big(k+1,\sqrt{k}\Big)\leq\frac{12}{k+1}\tan\theta_{c}\left(12,\,\frac{12}{k+1}\sqrt{k}\right)<\frac{12}{k+1}\tan\theta_{c}\left(12,\sqrt{12}\right)$
exists.
When $\alpha^{2}=\frac{k\alpha_{\tau}^{2}}{k_{\tau}}$ either for $\tau=1$ or for $\tau=2$,
(4.16) $\tan\theta_{c}\Big(k+1,\alpha\Big)$
$<$ $\frac{k_{\tau}+1}{k+1}\tan\theta_{c}\left(k_{\tau}+1,\,\frac{k_{\tau}+1}{k+1}\alpha\right)$
$=$ $\frac{k_{\tau}+1}{k+1}\tan\theta_{c}\left(k_{\tau}+1,\frac{\sqrt{k_{\tau}}+\frac{1}{\sqrt{k_{\tau}}}}{\sqrt{k}+\frac{1}{\sqrt{k}}}\alpha_{\tau}\right)$
$<$ $\frac{k_{\tau}+1}{k+1}\tan\theta_{c}\left(k_{\tau}+1,\alpha_{\tau}\right)$
and thus $\theta_{c}(k+1,\alpha)$ exists by the assumption $C(L_{\tau})$ is a Type-c cone.
Thus, no matter which value $\alpha^{2}$ takes from $\left\{k,k\frac{\alpha^{2}_{1}}{k_{1}},k\frac{\alpha^{2}_{2}}{k_{2}}\right\}$, we see from the above that $\theta_{c}(k+1,\alpha)$ always exists. ∎
We are ready to obtain more configuration results in this subsection.
Theorem 4.19.
Suppose $L_{1},L_{2}$ span Type-c cones with $k_{1},\,k_{2}\geq 3$. If $(L_{1})=s$$(L_{2})\geq 1$, then their minimal product $L_{1}\times L_{2}$ spans a Type-c cone of class (II).
Proof.
Without loss of generality, assume that $\cos R=1-\lambda_{1}^{2}(1-\cos R_{1})$. There are two cases to be discussed separately.
Case 1. $R_{1}\leq\frac{\pi}{2}$. In this case we can apply the key control Lemma 4.9 under the assumption on dimensions ($k_{1},\,k_{2}\geq 3$ would be enough for (4.3), as condition (2) of Lemma 4.9 satisfied) to get
$\tan\frac{R}{2}$ $\geq$ $\frac{k_{1}+1}{k_{1}+k_{2}+1}\tan\frac{R_{1}}{2}$
$\geq$ $\frac{k_{1}+1}{k_{1}+k_{2}+1}\tan\big(\theta_{c}(k_{1}+1,\alpha_{1})\big)$
$>$ $\tan\big(\theta_{c}(k_{1}+k_{2}+1,\frac{k_{1}+k_{2}+1}{k_{1}+1}\sqrt{{\tt s}(L_{1})k_{1}})\big)$
$=$ $\tan\big(\theta_{c}(k_{1}+k_{2}+1,\frac{\sqrt{k_{1}+k_{2}}+\frac{1}{\sqrt{k_{1}+k_{2}}}}{\sqrt{k_{1}}+\frac{1}{\sqrt{k_{1}}}}\sqrt{{\tt s}(L_{1})(k_{1}+k_{2})})\big)$
(4.18) $>$ $\tan\big(\theta_{c}(k_{1}+k_{2}+1,\alpha)\big)$
The last inequality (4.18) is due to the increasing property of $x+\frac{1}{x}$ when $x\geq 1$, the increasing property of $\theta_{c}(m,\cdot)$ in the second slot  [^5] This is clear in [Law] but to be self-contained we give a short proof in appendix., and the fact that $\alpha^{2}=(k_{1}+k_{2}){\tt s}(L_{1})$ under the assumption.
Case 2. $\frac{\pi}{2}<R_{1}\leq\pi$. By the proof of Lemma 4.18 for $(L_{1})=s$$(L_{2})\geq 1$ the vanishing angle $\theta_{c}(k+1,\alpha)$ always exists without bothering (4.15) so that we have no assumption on total dimension. Moreover, there is a free rough control $\theta_{c}(k+1,\alpha)\leq\tan\theta_{c}(k+1,\alpha)<\frac{1}{\alpha}$ since the function $(1-\alpha t)e^{\alpha t}$ vanishes at $t=\frac{1}{\alpha}$ (see appendix for more information about the equivalent calibration inequality where solution needs to reach zero before $t=\frac{1}{\alpha}$).
Under the assumption of Case 2, we have $\cos R\leq 1-\lambda_{1}^{2}$ which implies
$2\left(\frac{R}{2}\right)^{2}\geq 2\sin^{2}{\frac{R}{2}}=1-\cos R\geq\lambda_{1}^{2},$
and hence $\frac{R}{2}\geq\sqrt{\frac{k_{1}}{2k}}$. As $\alpha\geq\sqrt{k}$ , we also get $\frac{R}{2}\geq\theta_{c}(k+1,\alpha)$ when $k_{1}\geq 2$.
Therefore, in both Case 1 and Case 2, one can apply Lawlor’s criterion and the proof gets completed. ∎
Remark 4.20.
Examples of class (b) are many, see Example 4. All elements of class (b) has slope one. This theorem shows its power and derives a universal structure for the category of Type-c cones of class (b).
Not as in Theorem 1.1 where sufficiently many copies of a given $L$ are needed, for links of Type-c cones of class (II) two copies can work.
Corollary 4.21.
Let $L^{k}$ be the link of a Type-c cone of class (II) with $k\geq 3$ and $n\geq 2$. Then the minimal product $L^{\times n}$ can also span a Type-c cone of class (II).
From the proof of Theorem 4.19, we know that as long as $\cos R$ and $\alpha^{2}$ are determined by the same slot of the minimal product then the minimal product spans a Type-c cone. There are many specific cases for which this coincidence in the same slot occurs. A general structure result is the following (cf. Theorem 1.3).
Theorem 4.22.
Let $L^{k}$ be the link of a Type-c cone of class (II) and $n\geq k\geq 3$. Then $L\times\mathbb{S}^{n}$ can span a Type-c cone.
Proof.
Under the dimension assumption, we can see from Lemma 3.2 that
$\cos R=1-\frac{k}{n+k}(1-\cos R(L)).$
Moreover, by Lemma 3.1 we have $\alpha^{2}(L\times\mathbb{S}^{n})=(k+n)\max\{1,{\tt s}(L)\}=(k+n){\tt s}(L)$. Consequently, following the reasoning in the preceding paragraph, the conclusion stands true. ∎
Moreover, a subtle structure result related to Theorem 1.1 can be obtained.
Theorem 4.23.
Suppose $L=L_{1}^{\times s_{1}}\times\cdots\times L_{n}^{\times s_{n}}$ with $n\geq 2$, $s_{i}\geq 1$ and each component dimension $k_{i}>0$ for $i=1,\cdots,n$ spans a Type-c cone. Then so does $\tilde{L}\triangleq L_{1}^{\times(s_{1}+\ell_{1})}\times\cdots\times L_{n}^{\times(s_{n}+\ell_{n})}$ where $\ell_{i}\geq 0$ and $\sum_{i+1}^{n}k_{i}\ell_{i}\geq 3$.
Remark 4.24.
Here $L_{i}$ itself may not span area-minimizing cone.
Remark 4.25.
If we use $L\prec\check{L}$ to describe the relation of increasing exponential order between $L$ and $\check{L}$, then, given a strictly increasing sequence of this relation, according to Remark 4.1 there must be a (largest) successive infinite tail elements of which all span area-minimizing cones. The result here indicates that the tail and the element in the sequence which spans an area-minimizing cone of lowest dimension has gap at most $2$ in dimension.
Proof.
Without loss of generality, we consider $\check{L}\triangleq L_{1}^{\times(s_{1}+1)}\times\cdots\times L_{n}^{\times s_{n}}$ with $k_{1}\geq 3$ only and other situations can be derived accordingly. Note that by Lemma 4.6 it follows $R(L)\leq\frac{\pi}{2}$. By Proposition 4.13 or Lemma 3.1 we have ${\tt s}(L)\geq\max\{1,{\tt s}(L_{1})\}$. Therefore, both $R(\check{L})$ and ${\tt s}(\check{L})$ are decided by the first input $L$ of the minimal product of $L$ and $L_{1}$. When $k_{1}\geq 3$, the key control Lemma 4.9 applies and the argument in the proof in Case 1 of Theorem 4.19 leads us to the conclusion of the theorem. ∎
In fact more subtle structures can be established. To illustrate the point, let us mention two more situations, based on which interested readers may figure out many others.
Theorem 4.26.
Let $L^{k}$ be the link of a Type-c cone with $k\geq 3$. If $L$ is of class (II) then $L^{k}\times\mathbb{S}^{d_{1}}\times\cdots\mathbb{S}^{d_{n}}$ spans a Type-c cone.
Proof.
The reason is following. If some slot $\mathbb{S}^{d_{\tau}}$ decides $R$ instead of component $L$, then we are in Case 2 in the proof of Theorem 4.19 and can get $\frac{R}{2}\geq\sqrt{\frac{d_{\tau}}{d}}$ now and $\theta_{c}(d+1,\alpha)<\frac{1}{\alpha}\leq\sqrt{\frac{1}{d}}\leq\sqrt{\frac{d_{\tau}}{d}}$ by the existence of $\theta_{c}$.
When $R$ is decided by $L$, one only needs to note that $\alpha$ is also decided by $L$ and the statement follows according to the proof of Theorem 4.19. ∎
Instead of successively multiplying spheres, one can also consider using other kinds of links of class (I) (or their mixed types). For example, let $\tilde{M}$ be a focal submanifolds with ${\tt g}=4$ of sufficiently large dimension $n$. Since normal radius of each of them is always $\frac{\pi}{2}$, we get the following.
Theorem 4.27.
Let $L^{k}$ be the link of a Type-c cone of class (II) and $\frac{n}{2}\geq k\geq 3$. Then $L\times\tilde{M}^{n}$ can span a Type-c cone.
Proof.
By the same kind of reasoning in the proof of Theorem 4.22. ∎
Remark 4.28.
One can take $L$ to be a resulting Type (II) cone from Theorem 1.1. Then Theorems 4.22, 4.26 and 4.27 explore some extra structures. Moreover, if $L$ is of class (b), then so are the resultings in these theorems.
Remark 4.29.
We would like to point out that Lemma 4.18, Theorem 4.19, Remarks 4.20, Corollary 4.21, Theorems 4.22, 4.26 and 4.27 will also be true if “Type-c” is replaced by “Type-F”. The key point is to verify (4.16) and (4.18) with the extra factor from (2.5). We provide details in appendix.
Remark 4.30.
When “Type-c” is replaced by “Type-F”, the requirement the total dimension $\geq 11$ in Lemma 4.18 and Theorem 4.26 can be replaced by the total dimension $\geq 7$, since corresponding upper bound of replaced (4.15) exists as $\theta_{F}(8,\sqrt{7})<16^{\circ}$ according to Lawlor’s table in [Law]. Also see $(\ast)$ in appendix for remaining the requirement $k_{1},k_{2}\geq 3$” unchanged in the Type-F version of Theorem 4.19.
Based on the configuration results of this paper, in [z0] we get a constellation of area-minimizing cones of codimension larger than one which cannot be calibrated by any globally defined continuous calibrations.
5. Appendix
We collect proofs of Lemma 2.4, Footnote 5 and Remark 4.29 in appendix.
5.1. Proof of Lemma 2.4.
As briefly mentioned in §2 that Lawlor’s construction of area-decreasing projection is somehow equivalent to an inequality for the dual form to be a calibration. With $t=\tan\theta$ and $r$ the dimension of the cone, the inequality for calibration form is the following
$\left(g-\frac{tg^{\prime}}{r}\right)^{2}+\left(\frac{g^{\prime}}{r}\right)^{2}\leq\det\left(\textbf{I}-t\,\textbf{h}_{ij}^{v}\right)^{2},\text{ \ \ \ \ \ \ with }g(0)=1\ \ \text{ and \ \ }g\text{\ reaches }0.$
As Lawlor uses a uniform curve $\gamma$, the above inequality becomes
$\left(g-\frac{tg^{\prime}}{r}\right)^{2}+\left(\frac{g^{\prime}}{r}\right)^{2}\leq\inf_{x\in L,\,v\in S_{x}}\det\left(\textbf{I}-t\,\textbf{h}_{ij}^{v}\right)^{2}$
and being a Type-F cone means that there exists $g(t)$ such that
(5.1) $\left(g-\frac{tg^{\prime}}{r}\right)^{2}+\left(\frac{g^{\prime}}{r}\right)^{2}\leq\left(F(\alpha,t,r)\right)^{2},\text{ \ \ \ \ \ \ with }g(0)=1\ \ \text{ and \ \ }g\text{\ reaches }0.$
Recalling (2.3), we want to compare
(5.2) $F(\alpha,t,r)=\left(1-\alpha t\sqrt{\frac{r-1}{r}}\right)\left(1+\frac{\alpha t}{\sqrt{r(r-1)}}\right)^{r-1},$
and
(5.3) $F(\beta,s,m)=\left(1-\beta s\sqrt{\frac{m-1}{m}}\right)\left(1+\frac{\beta s}{\sqrt{m(m-1)}}\right)^{m-1}.$
Set $x=\alpha t\sqrt{\frac{r-1}{r}}=\beta s\sqrt{\frac{m-1}{m}}$ and they have simpler expressions
(5.4) $\left(1-x\right)\left(1+\frac{x}{{r-1}}\right)^{r-1}\text{\ \ \ and \ \ \ }\left(1-x\right)\left(1+\frac{x}{{m-1}}\right)^{m-1}.$
Note that, with fixed $x>0$,
(5.5) $\frac{d}{dn}\left(\log\left(1+\frac{x}{n}\right)^{n}\right)=\log\left(1+\frac{x}{n}\right)+n\cdot\frac{-\frac{x}{n^{2}}}{1+\frac{x}{n}}=\log\left(1+\frac{x}{n}\right)-\frac{x}{n+x}.$
Now taking derivative with respect to $x$ we have
$\frac{d}{dx}\left\{\log\left(1+\frac{x}{n}\right)-\frac{x}{n+x}\right\}=\frac{\frac{1}{n}}{1+\frac{x}{n}}+\frac{-n-x+x}{(n+x)^{2}}=\frac{1}{n+x}-\frac{n}{(n+x)^{2}}=\frac{x}{n+x}>0.$
As (5.5) is zero at $x=0$, we know (5.5) is positive when $x>0$. Hence the latter in (5.4) is larger than the former since $m>r$.
Inspired by the proof of Lemma 2.3 in [Law], we choose $\beta$ such that
$t=\frac{\beta}{\alpha}s\sqrt{\frac{(m-1)r}{m(r-1)}}=\frac{m}{r}s,$
namely,
$\beta\sqrt{\frac{(m-1)r}{m(r-1)}}=\frac{m}{r}{\alpha}.$
Set $h(s)=g(t)=g(\frac{m}{r}s)$. Then $\dot{h}=\frac{m}{r}g^{\prime}(\frac{m}{r}s)$. It follows from (5.1) for $g,r$ and $\alpha$ that
$\left(h(s)-\frac{s\dot{h}}{r}\right)^{2}+\left(\frac{\dot{h}}{m}\right)^{2}\leq\left(F(\alpha,t,r)\right)^{2}\leq\left(F(\beta,s,m)\right)^{2}.$
Since $\dot{h}\leq 0\leq h$, we have
$\left(h(s)-\frac{s\dot{h}}{m}\right)^{2}+\left(\frac{\dot{h}}{m}\right)^{2}\leq\left(F(\beta,s,m)\right)^{2}.$
This means that $h$ solves the calibration inequality and hence, by $t=\frac{m}{r}s$, we get
(5.6) $\tan\left(\theta_{F}\left(m,\frac{m}{r}\sqrt{\frac{(r-1)m}{(m-1)r}}\alpha\right)\right)<\frac{r}{m}\tan\left(\theta_{F}\left(r,\alpha\right)\right).$
To compare with (2.4) one can move the extra factor (inverse) to the right hand side and end up with Lemma 2.4.
5.2. Proof of Footnote 5.
To show Footnote 5 about the monotonicity property for $\theta_{c}(r,\cdot)$ in the second slot, one just needs to observe that $(1-\alpha t)e^{\alpha t}$ in (2.2) is decreasing in $\alpha$. So, suppose that that
$\left(h(s)-\frac{s\dot{h}}{r}\right)^{2}+\left(\frac{\dot{h}}{r}\right)^{2}\leq\left((1-\beta s)e^{\beta s}\right)^{2}$
holds. By setting $t=s$ and $g(t)=h(s)$, when $\beta>\alpha$ one has
$\left(g(t)-\frac{t\dot{g}}{r}\right)^{2}+\left(\frac{g^{\prime}}{r}\right)^{2}\leq\left((1-\beta s)e^{\beta s}\right)^{2}<\left((1-\alpha t)e^{\alpha t}\right)^{2}.$
Therefore, $\theta_{F}(r,\alpha)<\theta_{F}(r,\beta)$.
Remark 5.1.
The increasing monotonicity for $\theta_{F}(r,\cdot)$ follows similary.
5.3. Proof of Remark 4.29.
Note that to have (4.18) for Type-F we need to check
$\sqrt{\frac{(k_{1}+k_{2}+1)k_{1}}{(k_{1}+k_{2})(k_{1}+1)}}\frac{\sqrt{k_{1}+k_{2}}+1\big/\sqrt{k_{1}+k_{2}}}{\sqrt{k_{1}}+1\big/\sqrt{k_{1}}}>1$
which is equivalent to
$\sqrt{\frac{k_{1}+k_{2}+1}{k_{1}+1}}\frac{1+\frac{1}{k_{1}+k_{2}}}{1+\frac{1}{k_{1}}}>1.$
For this we check whether $\frac{(x+1)^{3}}{x^{2}}$ is increasing in $x$. By differentiating $3\log(x+1)-2\log x$, we have $\frac{3}{x+1}-\frac{2}{x}=\frac{x-2}{x(x+1)}>0$ when $x>2$. Hence the remark gets proved $(\ast)$ when $k_{1}\geq 2$ or when $k_{1}=1$ and $k_{2}\geq 4$.
Due to exactly the same reason (4.16) and with extra factor added also holds.